### v2

# Qingyu Shi1  Lu Qi2  Jianzong Wu1   Jinbin Bai3

Abstract
Customized image generation is crucial for delivering personalized content based on user-provided image prompts, aligning large-scale text-to-image diffusion models with individual needs. However, existing models often overlook the relationships between customized objects in generated images. Instead, this work addresses that gap by focusing on relation-aware customized image generation, which aims to preserve the identities from image prompts while maintaining the predicate relations described in text prompts. Specifically, we introduce RelationBooth, a framework that disentangles identity and relation learning through a well-curated dataset. Our training data consists of relation-specific images, independent object images containing identity information, and text prompts to guide relation generation. Then, we propose two key modules to tackle the two main challenges—generating accurate and natural relations, especially when significant pose adjustments are required, and avoiding object confusion in cases of overlap. First, we introduce a keypoint matching loss that effectively guides the model in adjusting object poses closely tied to their relationships. Second, we incorporate local features from the image prompts to better distinguish between objects, preventing confusion in overlapping cases. Extensive results on three benchmarks demonstrate the superiority of RelationBooth in generating precise relations while preserving object identities across a diverse set of objects and relations. The source code and trained models will be made available to the public.
Figure 1: Relation-aware image customization, the generated image should strictly keep the predicate relation and preserve each identity among those identities provided by the text prompt.
1 Introduction
Driven by large-scale text-to-image diffusion models (Rombach et al., 2022; Saharia et al., 2022; Podell et al., 2023; Pernias et al., 2023), customized image generation has recently made significant strides (Ruiz et al., 2023; Kumari et al., 2023; Wei et al., 2023; Zhang et al., 2024; Wang et al., 2024; Kong et al., 2024; Bai et al., 2024). This task focuses on generating images that preserve the identity of objects from user-provided inputs, enabling the creation of personalized and meaningful content. It has shown value in numerous applications, including personalized artwork, branding, virtual fashion try-ons, social media content creation, augmented reality experiences, and more.
Despite the success of many methods for customizing single or multiple objects (Liu et al., 2023; Ye et al., 2023; Li et al., 2023; Patel et al., 2024; Wang et al., 2024; Gu et al., 2024; Lin et al., 2024; Pang et al., 2024), they often overlook the relationships between objects and the corresponding text prompts. For instance, when two user-provided objects are paired with a text prompt specifying a particular relationship, the generated output should not only preserve their identities but also accurately reflect the intended relationship, such as a ‘hug’. This introduces new challenges in what we refer to as relation-aware customized image generation, which focuses on preserving multiple identities while adhering to relationship prompts.
An intuitive solution to this issue is to adapt existing tuning-based or training-based methods for customizing relationships (Huang et al., 2023; mengmeng Ge et al., 2024; Materzynska et al., 2023) among objects. However, both approaches face challenges. Tuning-based methods struggle to preserve multiple identities, as they invert objects into specific tokens, while training-based schemes often fail to balance image and text prompts, frequently overlooking key textual elements and hindering the generation of relationships between objects. As illustrated in Fig. 1, previous methods fail to capture the actions described in the text prompt and lose identity preservation.
We attribute this failure to two key factors: a lack of relevant data and an ineffective model design. Unlike data augmentation techniques such as flipping or rotation, commonly used to create paired training data in object customization methods (Yang et al., 2023; Chen et al., 2024b), our approach requires a triplet of images: two image prompts and one target image. The image prompts should contain similar objects but exhibit distinct actions compared to the target image. To collect these triplets, we propose a data engine to curate our fine-tuning set. We leverage an advanced text-to-image generation model (Betker et al., 2023) to generate triplets where the same object pair is shared across the images. Through text prompt guidance, the object in the image prompt provides strong identity information while maintaining distinct actions in the target image, enabling the model to focus on relation learning during fine-tuning.
For the model design, we propose RelationBooth, which applies the Low-Rank Adaptation (LoRA) (Hu et al., 2022) strategy to the text cross-attention layers of existing diffusion models to process user-provided text prompts. In RelationBooth, two key modules are introduced during training to enhance the customization of relationships and identity preservation. First, we introduce a keypoint matching loss (KML) as additional supervision to explicitly encourage the model to adjust object poses, since relationships between objects are closely tied to their poses. Importantly, the KML operates on the latent representation rather than the original image space, aligning with the default diffusion loss. Second, we inject local tokens for multiple objects to improve the distinctiveness of highly overlapping objects. Specifically, we employ the self-distillation method from CLIPSelf (Wu et al., 2024b) to enhance region-language alignment in CLIP’s dense features. Through partitioning and pooling, the well-aligned local tokens help mitigate appearance confusion between objects.
To more comprehensively evaluate relation-aware customized image generation, we constructed our RelationBench based on three established benchmarks (Ruiz et al., 2023; Kumari et al., 2023). Our methods demonstrate strong performance compared to existing approaches, achieving significant improvements in visual quality and quantitative results.
The contributions of this work are:
  • We explore a novel task called relation-aware customized image generation, which aims to preserve multiple identities from image prompts while adhering to the relationships specified in text prompts. This task can enhance various user-driven applications by enabling more control and customization.
  • We introduce a data engine that uses an advanced text-to-image generation model to generate triplet images where the same object pair is present with distinct actions. This well-curated dataset enables the model to focus on learning object relations during fine-tuning, leveraging the robust identity information provided by the image prompts.
  • Our proposed RelationBooth method has two key modules, including keypoint matching loss and object-related local tokens, to enhance the customization of relationships and identity preservation. The extensive experiments on three benchmarks demonstrate the effectiveness of our method.
2 Related Work
Diffusion-Based Text-to-Image Generation. This direction aims to produce high-quality images based on user-provided textual prompts. Diffusion-based generation models (Rombach et al., 2022; Saharia et al., 2022) encode these prompts through the text model (Radford et al., 2021; Chung et al., 2022) and inject the text embedding into the U-Net’s cross-attention layers. Some methods (Podell et al., 2023; Pernias et al., 2023; Ren et al., 2024; Wu et al., 2024a) upscale diffusion models and incorporate additional conditions as priors to generate high-resolution images. Meanwhile, several methods (Betker et al., 2023; Lian et al., 2023) introduce stronger text encoders or large language models (LLMs) to enhance the textual comprehension capabilities of diffusion models. Recent works (Peebles & Xie, 2023; Chen et al., 2024a) have replaced the U-Net denoiser with a Transformer-based denoiser. However, these models primarily focus on textual conditions and struggle to handle other forms of input, such as user-provided images. In contrast, our work facilitates customized generation while following predicate relation conditions given by text inputs.
Diffusion-Based Customized Generation. Customized image generation aims to produce diverse images based on user-provided concepts. Tuning-based methods accomplish this by fine-tuning specific parameters of diffusion models, thereby incorporating new concepts into text-to-image diffusion models like Stable Diffusion. Some of these methods (Gal et al., 2023; Voynov et al., 2023; Liu et al., 2023) employ text embeddings to represent customized concepts, which are then injected into prompts during inference. On the other hand, Dreambooth (Ruiz et al., 2023) fine-tunes the entire U-Net and introduces a prior preservation loss to mitigate language drift. Additionally, several approaches (Kumari et al., 2023; Gu et al., 2024; Lin et al., 2024; Pang et al., 2024) combine text embeddings and diffusion models for image synthesis. They fine-tune both the text embedding and the U-Net, yielding impressive results. Considering the substantial cost of fine-tuning for commercial uses, training-based methods (Li et al., 2023; Wei et al., 2023; Chen et al., 2024b; Patel et al., 2024; Zhang et al., 2024) have been proposed. This approach primarily utilizes an ID encoder to extract the concept’s identity and inject it into the U-Net via cross-attention layers. For instance, MS-Diffusion (Wang et al., 2024) integrates grounding tokens with a feature resampler to maintain detail fidelity and employs layout guidance to explicitly place the concepts. Despite their efficiency during inference, training-based methods often struggle with identity preservation and textual condition following ability. A typical problem is the relation expressions in prompts are often overlooked. To address this issue, we present RealtionBooth, to tackle relation-aware customized generation.
Relation-Aware Text-to-Image Generation. Inspired by tuning-based customization methods (Gal et al., 2023), some recent works (Huang et al., 2023; Wei et al., 2024; mengmeng Ge et al., 2024) represent a “neglected word” by learnable parameters. These methods fine-tune part of the parameters on content co-existing images. For instance, Reversion (Huang et al., 2023) fine-tunes the text embedding on a set of co-existing related images and introduces a relation-steering contrastive loss. Reversion gains a better alignment between relational words and generated images by injecting new text embedding into prompts. While effective with prepositions and adjectives, Reversion struggles with predicate relation which involves significant overlaps, and cannot customize user-provided concepts. Moreover, this method still struggles to generate vivid relationships while maintaining the fidelity of multiple customized concepts, likely due to the inherent disregard for text embeddings in customization methods. Our work bridges the gap in the predicate relation-following capability of training-based customization methods. This enables efficient inference that naturally and accurately generates predicate relations from the textual prompt.
3 Relation-aware customized obejct generation
Figure 2: The overview of RelationBooth. RelationBooth encodes the image prompts by dual encoder architecture. After getting the U-Net output $\hat{\epsilon}$, we predicate $\hat{z}_{0}$ and calculate the keypoint matching loss. The part in the dotted box is only for training.
3.1 Problem definition
Different from convention customization tasks, we explore a new setting that focuses on image generation by both image prompts $c_{i}\in\mathcal{R}^{N\times H\times W}$ and text prompts $c_{t}$. The $N$, $H$, and $W$ are the number, height, and width of image prompts. We call this setting RelationBooth due to the requirements that the generated image $\hat{x}$ should strictly preserve each identity and keep the predicate relation among those identities provided by $c_{i}$ and $c_{t}$.
We define this task as:
$\hat{x}=\Phi_{\theta}(c_{i},c_{t})$ (1)
where $\Phi$ is the network parametered by $\theta$. For brevity clarification, we set $N=2$ due to the basic triplet element defined in the RelationBooth.
3.2 Discussion and Motivation of RelationBooth
Although previous works (Kumari et al., 2023; Gu et al., 2024; Zhang et al., 2024) have attempted to customize multiple objects under text control, we are the first to address the generation of relation-aware customized images. Notably, the state-of-the-art multi-object customization model, MS-Diffusion (Wang et al., 2024), cannot handle this task effectively. Specifically, given the image prompt $c_{i}$ and text prompt $c_{t}$. MS-Diffusion utilizes the CLIP (Radford et al., 2021) to extract the image and text tokens from $c_{i}$ and $c_{t}$ respectively. The image and text tokens are injected into the U-Net $\bm{\epsilon_{\theta}}$ through parallel cross-attention layers. The U-Net is optimized on the latent representation of image $z$:
$\mathcal{L}_{MS}=\mathbb{E}_{z_{t},t,\epsilon,c}[\|\epsilon_{t}-\bm{\epsilon_{\theta}}(z_{t},t,c_{i},c_{t})\|_{2}^{2}]$ (2)
where $\epsilon_{t}\sim\mathcal{N}(0,1)$, we omit the CLIP encoder for brevity. However, despite MS-Diffusion can customize objects under text control, it’s struggled for MS-Diffusion to generate relations accurately. Furthermore, MS-Diffusion might cause object confusion when heavy overlapping exists, leading to failure relation generation.
We revise the relation generation task. Existing relation inversion methods (Huang et al., 2023; mengmeng Ge et al., 2024) aim to invert a relation to a text embedding $R^{*}$ in the pre-trained text-to-image diffusion models. Given a set of relation-specific images $\{x_{k}\}_{k=1}^{n}$, they employ denoising loss $\mathcal{L}_{RI}$ to fine-tune $R^{*}$ for alignment with specific relation:
$\mathcal{L}_{RI}=\mathbb{E}_{z_{t},t,\epsilon,c}[\|\epsilon_{t}-\bm{\epsilon_{\theta}}(z_{t},t,c_{t})\|_{2}^{2}]$ (3)
Directly implementing relation inversion methods on MS-Diffusion leads to suboptimal results. We fine-tune MS-Diffusion on a set of relation-specific images, and experiment results show the model falls short in generating relation in Fig. 4. We think the image prompt $c_{i}$ that is not considered in relation inversion methods affects the generation process, leading to an overlook of $R^{*}$. In the following subsections, we will introduce the data collection and proposed method, considering lacking both of them in the proposed tasks.
3.3 Data Collection
Considering the scarcity of available datasets that can be used, we first design a data engine for collecting high-quality tuning data. We denote the relation-specific image as $x$. While $x$ contains relation information, it also includes the object’s identities. Directly fine-tuning on a set of $x$ leads to unsatisfied relation generation. To disentangle the relation and identity in $x$. The ideal tuning data should be in the form of $\mathcal{D}=(x,c_{i},c_{t})$. Where $c_{i}$ represents the image prompts that contain the object in $x$, $c_{t}$ represents the text prompt that guides the relation generation. Unfortunately, directly cropping the objects from $x$ to obtain $c_{i}$ results in a copy-and-paste effect, even when using data augmentation techniques like flipping and rotation. Therefore, we introduce our data engine below.
Data engine. Fig. 3 illustrates our data generation process. Inspired by the capability of DALL-E 3 and its multi-turn dialog capability. We use DALL-E 3 to generate $\mathcal{D}=(x,c_{i},c_{t})$. Where $x$ and $c_{i}$ share the same object identity by leveraging the prompt “The photo of the same.” We observe that prompting in this manner allows DALL-E 3 to remember and preserve identity in common categories such as “tiger” and “brown bear” which are enough for relation learning. Next, we employ X-Pose (Yang et al., 2024), SAM (Kirillov et al., 2023), and LLaVA (Liu et al., 2024) to annotate $x$, $c_{i}$ with keypoints, masks, and captions.
Figure 3: The overview of the data generation process. DALLE-3 can generate similar objects by using ’The photo of the same’ in the text prompt on common animal categories.
Figure 4: Comparison of different fine-tuning data. Simply implementing relation inversion methods on MS-Diffusion leads to failure relation generation.
3.4 RelationBooth
3.4.1 RelationLoRA Fine-tuning
To disentangle the relation and identity information in relation-specific images $x$, we utilize off-the-shelf identity inject methods from state-of-the-art customized model MS-Diffusion (Wang et al., 2024). Specifically, we utilize parallel cross-attention layers to process the $c_{i}$ and $c_{t}$, the identity information in $c_{i}$ facilitates the model to focus on relation information in $x$ and $c_{t}$ during fine-tuning. We formulate the parallel cross-attention layers:
$h=\gamma\cdot\operatorname{Softmax}(\frac{QK_{i}}{\sqrt{d}})V_{i}+\operatorname{Softmax}(\frac{QK_{t}}{\sqrt{d}})V_{t}$ (4)
where $Q=W_{q}h$, $K_{i}=W_{k_{i}}c_{i}$, $V_{i}=W_{v_{i}}c_{i}$, $K_{t}=W_{k_{t}}c_{t}$, $V_{t}=W_{v_{t}}c_{t}$, and $\gamma$ is a hyperparameter for scaling features from $c_{i}$. For clarity, we omit the final linear layer $W_{out}$. The predicate relation information mainly depends on text prompts. Therefore, we inject LoRA layers in all of the $W_{q}$, $W_{k_{t}}$, $W_{v_{t}}$, and $W_{out}$ to encourage the model to pay more attention to the predicate relation in $c_{t}$. We freeze all other parameters during fine-tuning. The required additional storage space for LoRA weights is only 13MB. As shown in Fig.4, benefited from the disentangled identity information in $c_{i}$, the LoRA weights effectively capture the predicate relation in $x$ and $c_{t}$, accurately generating the corresponding images. Notably, after fine-tuning, our RelationLoRA can also be directly integrated into SDXL (Podell et al., 2023), effectively enabling the model to address the relation inversion task. The extensive results are presented in the appendix (Fig. 15).
3.4.2 Keypoint Matching Loss
Most relations have specific requirements for the pose, such as “hugging” requires the arm to crossed over the other object’s body, and “riding a bicycle” requires the feet of the object to be on the pedal plate. Therefore, it’s important to accurately manipulate the pose of objects in the right status. Intuitively, we propose to introduce keypoint features into the fine-tuning procedure to explicitly guide the pose of the objects for better relation generation.
Specifically, considering our task involves common objects instead of humans only, We employ X-Pose (Yang et al., 2024) as keypoint detector because it can detect any keypoints in complex real-world scenarios. We detect 17 keypoints for each object in $x$ and $c_{i}$, and denote the keypoint coordinates as $c^{x}_{kp},c^{c_{i}}_{kp}\in\mathcal{R}^{17\times 2}$ respectively. To encourage the model to generate accurate pose in $\mathcal{D}(\hat{z}_{0})$, where $\mathcal{D}$ is the VAE decoder. We use the U-Net’s output $\hat{\epsilon}$ to predict $z_{0}$ during fine-tuning:
$\hat{z}_{0}=\frac{z_{t}-\sqrt{1-\bar{\alpha}_{t}}\hat{\epsilon}}{\sqrt{\bar{\alpha}_{t}}}$ (5)
We use the VAE encoder $\mathcal{E}$ to obtain latent representation of image prompts $\mathcal{E}(c_{i})$. Then we calculate the MSE loss on the corresponding keypoint locations between $\mathcal{E}(c_{i})$ and $\hat{z}_{0}$:
$\mathcal{L}_{KML}$ $=\frac{1}{n_{kp}}\mathbb{E}_{\bm{z_{t}},c_{i}}\left\|\mathcal{E}(c_{i})[c^{c_{i}}_{kp}]-\hat{z_{0}}[c^{z_{0}}_{kp}]\right\|^{2}_{2}$ (6)
$\mathcal{L}$ $=\mathcal{L}_{denoise}+\lambda\cdot\mathcal{L}_{KML}$ (7)
where $n_{kp}$ denote the number of keypoint, $c^{c_{i}}_{kp}$ and $c^{z_{0}}_{kp}$ are keypoint’s coordinates in $c_{i}$ and $z_{0}$ respectively, $\lambda$ controls for the relative weight of KML. The part inside the dotted box in Fig. 2 illustrates the model fine-tuning with the KML. We find KML is effective in encouraging the model to manipulate the pose for relation generation.
3.4.3 Local Token Injection
When generating relations, overlapping objects often occur. It is challenging to distinguish multiple objects solely using CLIP’s image-level embedding of the full image, which can lead to confusion between customized objects during relation generation. To address this limitation, we introduce a local region representation of images. Considering the domain shift from the full image to local image regions, we employ CLIPSelf’s self-distillation methods (Wu et al., 2024b) to enhance the region-language alignment of local region representation. Specifically, in line with CLIPSelf, we extract dense features (Zhou et al., 2022) from a student CLIP model and derive local tokens through partitioning and pooling (He et al., 2017). We partition the full image into patches and feed them into a frozen CLIP model to obtain teacher tokens. Our approach minimizes the cosine similarity between local tokens and teacher tokens to improve the region-language alignment of local tokens. To formulate this process, we denote the hidden states of the second-to-last layer of CLIP as $h^{\prime}_{clip}$. We modify the last layer of CLIP $h_{clip}=\operatorname{ModifiedAttention}(h_{clip})$ to obtain the dense features:
$h_{tmp}=\operatorname{Proj}_{v}(\operatorname{norm}(h^{\prime}_{clip})),h_{tmp}=h^{\prime}_{clip}+\operatorname{Proj}_{out}(h_{tmp}),h_{tmp}=h_{tmp}+\operatorname{FFN}(h_{tmp})$ (8)
where $\operatorname{Proj}$ means the linear projection. We discard the class embedding $h_{clip}[0]$ and reshape the image embeddings $z^{\prime}[1:h\times w]$ into an $h\times w$ feature map. We partition the feature map into patches, from which we can get local tokens $tok_{l}$ by pooling. During inference, we simply concatenate the aligned local tokens with the image-level tokens and send them to the ID extractor which is a transformer-based architecture shown in Fig. 2 b). The query $q$ is obtained by concatenating class text embedding with the bounding box.
$q=q+\operatorname{GatedSelfAttn}([q,tok_{i},tok_{l}])$ (9)
Experimental results show that local token injection enhances identity preservation in evaluation metrics and visual effects as shown in Fig. 9.
4 Experiments
Figure 5: Single-object comparison with our base model MS-Diffusion: The results demonstrate that our method generates more accurate and natural relation-aware images.
Table 1: Quantitative comparison on RelationBench, Bold and underline represent the highest and second-highest metrics.
Method Single-object Multi-object
CLIP-T CLIP-R CLIP-I DINO CLIP-T CLIP-R CLIP-I DINO
DreamBooth (SDXL) 27.8 18.2 74.2 62.6 24.3 16.2 67.8 57.2
Custom Diffusion 26.5 15.2 73.2 58.8 20.1 15.4 64.7 55.3
Cones-V2 24.4 13.5 72.1 57.2 21.3 15.2 64.3 54.2
ELITE 25.7 14.9 75.4 61.5
AnyDoor 24.5 14.7 77.4 62.2 21.6 14.9 69.7 59.8
BLIP-Diffusion 26.2 15.7 77.4 57.7
SSR-Encoder 25.5 15.9 80.4 59.4 24.2 14.6 72.1 56.2
MS-Diffusion 26.5 18.8 78.7 64.5 26.9 18.9 73.8 58.8
Ours 30.6 21.4 77.9 63.4 28.9 20.4 75.4 62.1
Figure 6: Multi-object comparison with training-based and tuning-based methods: Compared to our base model, our method shows a clear advantage in relation-aware generation and avoiding object confusion in overlapping scenarios.
Table 2: DreamBench.
Method Single-object
CLIP-T CLIP-I DINO
DreamBooth (SDXL) 31.2 81.5 69.2
Custom Diffusion 28.4 77.2 66.8
Cones-V2 31.0 76.5 67.2
ELITE 29.8 77.4 62.5
AnyDoor 25.5 82.1 67.8
SSR-Encoder 30.8 82.1 61.2
MS-Diffusion 31.5 79.3 68.2
Ours 31.6 78.9 67.4
Table 3: M-CustomConcept101.
Method Multi-object
CLIP-T CLIP-I DINO
DreamBooth (SDXL) 29.5 67.4 49.2
Custom Diffusion 28.1 66.2 48.8
Cones-V2 29.2 66.5 47.2
AnyDoor 20.2 72.1 51.2
SSR-Encoder 30.6 71.1 52.2
$\lambda$-eclipse 29.2 68.2 48.2
MS-Diffusion 28.0 70.2 51.2
Ours 29.4 71.4 52.3
Figure 7: Additional results of relation-aware generation.
Figure 8: Additional results of relation-aware generation across a wide range of objects.
Figure 9: Ablation study of keypoint matching loss (KML) and local tokens: The images are generated using the same random seed. We use green and red boxes to highlight the main differences. The areas within the green boxes show the improvements with our proposed component.
Table 4: Ablation study on our proposed components
Method Single-object Multi-object
CLIP-T CLIP-R CLIP-I DINO CLIP-T CLIP-R CLIP-I DINO
w/o Relation-aware Data 28.4 16.9 78.2 64.4 26.2 16.8 75.3 59.7
w/o Local Token Injection 29.2 21.0 76.8 62.5 28.5 19.5 75.1 59.3
w/o Keypoint Matching Loss 28.3 17.9 77.8 62.1 27.4 18.2 75.2 61.2
Full Model 30.6 21.4 77.9 63.4 28.9 20.4 75.4 62.1
Table 5: Ablation study on Local Image Encoder’s architecture.
Model Multi-object
CLIP-T CLIP-R CLIP-I DINO
EVA-CLIP-L14 23.6 15.7 56.4 54.8
CLIP-ViT-L14 22.9 14.8 58.3 52.7
CLIP-ViT-bigG 28.9 20.4 75.4 62.1
Table 6: Ablation study on $\lambda$, controlling relative weight of KML.
Lambda Multi-object
CLIP-T CLIP-R CLIP-I DINO
$\lambda$=1e-2 27.5 19.3 72.6 59.6
$\lambda$=1e-3 28.9 20.4 75.4 62.1
$\lambda$=1e-4 26.8 18.1 73.9 60.4
Table 7: Ablation study on the number of local tokens (determined by partition size).
Num Local Tokens Multi-object
CLIP-T CLIP-R CLIP-I DINO
N=2$\times$2 28.3 19.4 75.3 60.1
N=4$\times$4 28.9 20.4 75.4 62.1
N=8$\times$8 27.9 18.1 74.1 59.4
Table 8: Ablation study on local token injection methods to ID extractor.
Injection method Multi-object
CLIP-T CLIP-R CLIP-I DINO
Add 25.4 18.5 71.0 56.9
Linear Projection 25.8 18.3 68.2 54.4
Concatenate 28.9 20.4 75.4 62.1
Implementation Details. For relation-aware fine-tuning, we generate a tuning set with 4-6 samples per relation. We incorporate LoRA layers into all text cross-attention layers of the U-Net. We set the LoRA rank to $r=4$, the parallel cross-attention scaling factor to $\gamma=0.6$, and the keypoint matching loss weight to $\lambda=1e-3$. The model is fine-tuned for 500 steps, using 2 A100 GPUs, with a total batch size of 8, completing the process in 10 minutes. We use the Adam optimizer with a learning rate of $41e-$ 4 and no weight decay, resulting in a total of 3.1M trainable parameters. Note that our RelationLoRA only modifies the text cross-attention layers in SDXL, making it compatible with any SDXL-based models. Considering the strong identity preservation capabilities of MS-Diffusion, we implement our RelationBooth on MS-Diffusion. During fine-tuning and inference, we concatenate local tokens with image tokens. We put more self-distillation details of the Local Image Encoder in appendix (6.1). Additionally, as shown in appendix (Fig. 15), our RelationLoRA is compatible with SDXL for addressing the Relation Inversion task (Huang et al., 2023).
Evaluation. To evaluate relation-aware customized image generation, we propose RelationBench, consisting of 44 objects from DreamBench (Ruiz et al., 2023) and CustomConcept101 (Kumari et al., 2023), along with 25 predicate relations. The object categories include pets, plushies, toys, people, and cartoons. Using GPT-4, we generate 100 cases for single- and multi-object evaluations. Following previous works (Ruiz et al., 2023; Wang et al., 2024), we evaluate our method on three metrics: (1) Identity Preservation, which assesses the similarity between the generated images and the image prompts, using the CLIP image score and DINO score that calculate the cosine similarity between the class embeddings of images, referred to as CLIP-I and DINO, respectively; (2) Text Alignment, which evaluates how well the generated images align with the text prompts, using the CLIP image-text score, denoted as CLIP-T; and (3) Relation Alignment. We observed that nouns in the prompts can inflate CLIP-T scores, making them less accurate for evaluating relation generation. To address this, we extract relation predicate from the prompts using spaCy and calculate the CLIP image-text score using only the predicate, denoted as CLIP-R.
4.1 Main Results
We evaluate our method through both quantitative and qualitative results in single- and multi-object cases. In single-object cases, we primarily focus on the alignment of the object’s pose with the predicate relation in the text prompt, as well as the model’s ability to preserve identity. In multi-object cases, in addition to relation generation and identity preservation, we also need to ensure there is no confusion between the multiple objects.
Quantitative results. We quantitatively compare our method with baseline models across three benchmarks: RelationBench, DreamBench, and multi-object cases in CustomConcept101. As shown in Tab. 1, our method demonstrates a clear advantage on RelationBench for single-object cases, particularly in the CLIP-T and CLIP-R metrics. However, it performs slightly lower in the CLIP-I metric, likely due to significant pose variations in the objects, which can negatively impact the CLIP-I score. For multi-object generation, our method surpasses other approaches across all evaluation metrics, which we attribute to effectively preventing object confusion, leading to a significant improvement in CLIP-I performance. Furthermore, as shown in Tab. 3, we evaluate single-object generation performance on DreamBench. Our method outperforms others in the CLIP-T metric while delivering competitive results in CLIP-I and DINO. To evaluate multi-object generation, we conduct experiments on multi-object cases from CustomConcept101, which we denote as M-CustomConcept101. As shown in Fig. 3, our model achieves the highest performance in the DINO score and ranks second in both the CLIP-T and CLIP-I metrics.
Qualitative Comparison. We conduct qualitative comparison experiments for relation-aware customized image generation on RelationBench. For single-object comparisons, as shown in Fig. 5, our generated images are more accurate than those produced by our base model, MS-Diffusion, and naturally reflect the predicate relations in the text prompts. Additional single-object comparison results are provided in the appendix (Fig. 12). In multi-object comparisons, Fig. 6 demonstrates that our method excels in generating multi-object relations, successfully avoiding object confusion while accurately representing the intended relations.
4.2 Ablation study and Analysis
RelationLoRA Fine-tuning. Instead of injecting LoRA layers and fine-tuning on our curated $(x,c_{i},c_{t})$, we fine-tune the text embedding $R^{*}$ using a set of relation-specific data $(x,c_{t})$, where blank images are used as image prompts $c_{i}$. The visual comparison is presented in Fig. 4, and the quantitative results in Tab. 4 indicate that this method underperforms in both CLIP-T and CLIP-R scores. This may be due to the entanglement of identity and relation information in $x$, as well as the gap between the blank and real images, making it difficult to accurately generate the relation.
Keypoint Matching Loss. For the keypoint matching loss (KML), we omit it during fine-tuning. As shown in Fig. 9 a), including the KML leads to more accurate and natural relation-aware images when using the same random seed. Quantitative results, presented in Tab. 4, show a decline across all evaluation metrics when the KML is removed. Additionally, we experimented with different values for $\lambda$ and found that the best performance occurs at $\lambda=1e-3$.
Local Token Injection. For local token injection, the visual comparison in Fig. 9 b) shows severe object confusion when local tokens are omitted. The quantitative results in Tab. 8 demonstrate that local tokens improve relation-aware customization. We conduct further ablation as shown in Tab. 8, Tab. 6, and Tab. 8. We experimented with varying numbers of local tokens injected into the ID extractor and found that 16 tokens yielded the best results across all evaluation metrics. Therefore, we use 16 local tokens in all experiments. Additionally, we swapped the Local Image Encoder and found that CLIP-ViT-bigG was the most compatible model. For the injection method, we note that simply concatenating local tokens with image tokens produced the best outcomes. Combining these methods ensures that our method maintains high quality relation-aware customized generation.
5 Conclusion
In this work, we introduce a challenging task: relation-aware customized image generation, which aims to generate objects that adhere to the relations specified in the text prompt while preserving the identities of user-provided images. To support this, we propose a data engine for generating high-quality fine-tuning data. Our method, RelationBooth, incorporating Keypoint Matching Loss and Local Token Injection, effectively captures relation information and generates natural relations between customized objects while mitigating object confusion. We also provide a fair comparison with other methods on a new benchmark, RelationBench. The extensive experiments on three benchmarks demonstrate the effectiveness of our method, highlighting its potential for interactive scenario generation, relation detection, etc.
6 Appendix
This appendix provides additional details on the implementation of the Local Image Encoder and baseline methods. In the second section, we offer a brief introduction to the RelationBench benchmark. We then present further experimental results to validate the effectiveness of our method across various scenarios. Following that, we provide an in-depth analysis of the proposed components’ contributions to overall performance. Lastly, we discuss the potential ethical, privacy, and fairness implications of this technology, highlighting its social responsibility in real-world applications.
6.1 More Implementation Details
Local Image Encoder’s Implementation Details. To enhance region-language alignment of dense features, we employ self-distillation on CLIP-ViT-bigG. The training is conducted on the train2017 split of the COCO dataset for six epochs, using 8 A100 GPUs with a batch size of 2 per GPU. We apply the Adam optimizer with a learning rate of $51e-$ 5 and a weight decay of 0.1. The Local Image Encoder, containing 1.8B parameters, extracts local tokens that are concatenated with image tokens during fine-tuning and inference to improve the alignment of object appearance. This distillation process enhances the performance of our base model, MS-Diffusion, in multi-object generation tasks. However, significant object confusion still occurs during relation generation.
Baselines’ Implementation Details. For tuning-based methods, we implement Textual Inversion, DreamBooth, and Custom Diffusion using their respective Diffusers versions, with learning rates and tuning steps aligned to those reported in the original papers. We utilize the official implementations and checkpoints for training-based methods, adjusting hyperparameters as needed during evaluation. Specifically, we set the scale to 0.6 in MS-Diffusion and sample 30 steps using the EulerDiscreteScheduler. For the SSR Encoder, we employ the UniPCMultistepScheduler, sampling 30 steps and adjusting the scale for each object to accommodate different cases. For $\lambda$-Eclipse, we apply the default settings of the official implementation without modification.
6.2 RelationBench
In this section, we show the object and text prompt contained in our RelationBench in Fig. 10.
Figure 10: Objects in our proposed RelationBench
Figure 11: Object category in RelationBench
Table 9: Text prompt in RelationBench.
No. Prompt
1 A { } is playing guitar on a park bench, serenading passersby.
2 A { } is playing piano in a grand hall filled with glowing chandeliers.
3 A { } is eating dinner in a bustling restaurant with soft jazz playing in the background.
4 A { } is dancing in the moonlight on a rooftop terrace with a stunning city view.
5 A { } is lifting weights in a modern gym, pushing through an intense workout.
6 A { } is reading a book by the fireplace on a rainy evening.
7 A { } is skiing down a steep slope in the Alps, with snowflakes falling gently.
8 A { } is sleeping peacefully in a hammock under the shade of a palm tree.
9 A { } is cooking lunch in an open-air kitchen overlooking a lush green valley.
10 A { } is singing on stage during a vibrant music festival with thousands of fans.
11 A { } is riding a bike along the scenic countryside road during sunset.
12 A { } is riding a horse through a dense forest, surrounded by nature.
13 A { } is riding a motorbike on a winding mountain road, feeling the wind rush by.
14 A { } is playing soccer on a sandy beach, with a vibrant sunset in the background.
15 A { } is playing chess with a { } in a quiet park under the shade of a tree.
16 A { } is partner dancing with a { } in a vintage ballroom with live music playing.
17 A { } is carrying a { } across a rushing river, carefully finding their steps.
18 A { } is fencing with a { } in an elegant arena, with spectators watching closely.
19 A { } shakes hands with a { } at a formal awards ceremony after receiving a prize.
20 A { } is kissing a { } in the rain under a large umbrella in a romantic city square.
21 A { } is playing basketball with a { } on a street court.
22 A { } is wrestling with a { } in a championship ring.
23 A { } is hugging a { } in front of the mountain.
24 A { } is fighting with a { } in a garden.
25 A { } is sitting back to back with a { } on a hilltop.
6.3 More Results of Relation-aware Image Customization
More Results of Single-object relation-aware generation. We show more qualitative results in Fig. 12. We compare our methods with both training-based and tuning-based methods.
Figure 12: Single-object comparison. TI and DB indicate Textual Inversion and DreamBooth, respectively. Our methods achieve the best balance between relation generation and identity preservation.
Figure 13: Single-object comparison with our base model MS-Diffusion
More Results of Multi-object relation-aware generation. In this section, we present additional results on multi-object relation-aware generation. In addition to implementing our RelationBooth on MS-Diffusion, we also use SDXL as a base model. As shown in Fig. 15, our RelationBooth enhances SDXL’s ability to generate images that accurately depict object relations. Compared to ReVersion (Huang et al., 2023), our method generates more precise relations, without object confusion or missing objects, demonstrating the strong performance of RelationBooth in the Relation Generation task.
Figure 14: Multi-object relation-aware image customization results of pet, toy, plushie, and person.
Figure 15: Our RelationBooth is compatible with SDXL to address Relation Inversion task.
Failure cases. As shown in Fig. 16, we present three typical failure cases from our experimental results. The first involves unreasonable relation generation requests, such as asking a plushie octopus, which lacks limbs, to ’shake hands.’ In response, our model generates additional arms for the plushie octopus, leading to a mismatched appearance. The second issue is the unnaturalness of some generated relations, such as a duck failing to make contact with a cat as it should. The final failure case is object confusion at the interaction point, which is a common challenge across all multi-object generation models.
Figure 16: Failure cases of our RelationBooth.
6.4 Additional Analysis on Our Proposed Components
Our well-curated Data. Our curated data are generated by DALLE-3 (Betker et al., 2023), which can effectively preserve the identity of common categories, such as animals, through its multi-turn dialogue capability. While our fine-tuning dataset consists solely of animals, our method focuses primarily on the relation information in the ground truth by disentangling identity using an off-the-shelf identity extractor. Therefore, the specific categories in the fine-tuning data are not the most critical factor for relation learning.
Figure 17: Our fine-tuning dataset as an example.
Keypoint Matching Loss. We use X-Pose as our keypoint detector due to its open-vocabulary detection capabilities. The keypoint matching loss (KML) facilitates relation generation by explicitly guiding the model’s pose manipulation, resulting in more accurate pose generation. As shown in Fig. 19, the cat’s arm crosses the dog’s body, accurately depicting the ’hug’ relation.
Figure 18: Illustration of Keypoint Matching Loss
Figure 19: Additional ablation studies on KML and Local Tokens.
Local Tokens Visualization. To understand why local features enhance relation-aware generation, we adopt Principal Component Analysis (PCA) to project the dense feature more compactly. As shown in Fig. 20, dense features provide more fine-grained information than CLIP image tokens, aiding in distinguishing between different objects during the generation process and helping to avoid object confusion, especially in cases of heavy overlap. Moreover, it can facilitate object appearance alignment.
Figure 20: Visualization of Dense feature and Image tokens by Principal Component Analysis (PCA).
6.5 Social Compact
Positive societal impacts. Relation-aware image customization enables users to generate images that not only contain customized objects but also capture their meaningful relationships. This opens up new opportunities for creative professionals, such as designers, advertisers, and educators, to communicate complex ideas visually with greater precision and flexibility. It has the potential to streamline content creation in diverse fields, from personalized marketing to educational tools, making high-quality, contextually rich imagery accessible without the need for extensive resources.
Potential negative societal impacts. The ability to generate customized images that involve specific relationships between objects could be misused to fabricate misleading or harmful visual narratives, including false representations of events or manipulative visual content in political or social contexts. Additionally, if the models are trained on biased data, they may reinforce existing societal biases, marginalizing certain groups or distorting the accuracy of represented relationships.
Mitigation strategies. To reduce misuse, ethical guidelines should be established to govern the responsible development and application of this technology. Promoting transparency about generated content and integrating fairness and diversity considerations into dataset selection are key strategies for mitigating potential harms.
6.6 Incorporate with CogVideoX-5b-I2V
Figure 21: Incorporate with CogVideoX-5b-I2V.

### v3

@@
-Qingyu Shi1  Lu Qi2  Jianzong Wu1   Jinbin Bai3
+DreamRelation: Bridging Customization and Relation Generation
 Abstract
-Customized image generation is crucial for delivering personalized content based on user-provided image prompts, aligning large-scale text-to-image diffusion models with individual needs.
-However, existing models often overlook the relationships between customized objects in generated images.
-Instead, this work addresses that gap by focusing on relation-aware customized image generation, which aims to preserve the identities from image prompts while maintaining the predicate relations described in text prompts.
-Specifically, we introduce RelationBooth, a framework that disentangles identity and relation learning through a well-curated dataset.
+Customized image generation is essential for creating personalized content based on user prompts, allowing large-scale text-to-image diffusion models to more effectively meet individual needs.
+However, existing models often neglect the relationships between customized objects in generated images.
+In contrast, this work addresses this gap by focusing on relation-aware customized image generation, which seeks to preserve the identities from image prompts while maintaining the relationship specified in text prompts.
+Specifically, we introduce DreamRelation, a framework that disentangles identity and relation learning using a carefully curated dataset.
 Our training data consists of relation-specific images, independent object images containing identity information, and text prompts to guide relation generation.
-Then, we propose two key modules to tackle the two main challenges—generating accurate and natural relations, especially when significant pose adjustments are required, and avoiding object confusion in cases of overlap.
+Then, we propose two key modules to tackle the two main challenges—generating accurate and natural relationships, especially when significant pose adjustments are required, and avoiding object confusion in cases of overlap.
 First, we introduce a keypoint matching loss that effectively guides the model in adjusting object poses closely tied to their relationships.
-Second, we incorporate local features from the image prompts to better distinguish between objects, preventing confusion in overlapping cases.
-Extensive results on three benchmarks demonstrate the superiority of RelationBooth in generating precise relations while preserving object identities across a diverse set of objects and relations.
+Second, we incorporate local features of the image prompts to better distinguish between objects, preventing confusion in overlapping cases.
+Extensive results on our proposed benchmarks demonstrate the superiority of DreamRelation in generating precise relations while preserving object identities across a diverse set of objects and relationships.
 The source code and trained models will be made available to the public.
-Figure 1: Relation-aware image customization, the generated image should strictly keep the predicate relation and preserve each identity among those identities provided by the text prompt.
+11footnotetext: * Corresponding authors: Lu Qi, Xiangtai Li.
+Figure 1: In our Relation-Aware Image Customization task, the generated images must accurately preserve the relationships between objects and maintain their identity.
+We highlight the limitations of previous approaches using three color codes: red indicates failure to capture relationships, blue marks missing objects, and orange represents object confusion.
+Each image is annotated to reflect its specific issue.
+Our results, highlighted by green boxes, demonstrate the advantages of our proposed method.
 1 Introduction
-Driven by large-scale text-to-image diffusion models (Rombach et al., 2022; Saharia et al., 2022; Podell et al., 2023; Pernias et al., 2023), customized image generation has recently made significant strides (Ruiz et al., 2023; Kumari et al., 2023; Wei et al., 2023; Zhang et al., 2024; Wang et al., 2024; Kong et al., 2024; Bai et al., 2024).
+Driven by large-scale text-to-image diffusion models [30, 32, 27, 26], customized image generation has recently made significant strides [31, 15, 36, 43, 35, 14, 1].
 This task focuses on generating images that preserve the identity of objects from user-provided inputs, enabling the creation of personalized and meaningful content.
 It has shown value in numerous applications, including personalized artwork, branding, virtual fashion try-ons, social media content creation, augmented reality experiences, and more.
-Despite the success of many methods for customizing single or multiple objects (Liu et al., 2023; Ye et al., 2023; Li et al., 2023; Patel et al., 2024; Wang et al., 2024; Gu et al., 2024; Lin et al., 2024; Pang et al., 2024), they often overlook the relationships between objects and the corresponding text prompts.
+Despite the success of many methods for customizing single or multiple objects [20, 42, 16, 24, 35, 8, 18, 23], they often overlook the relationships between objects and the corresponding text prompts.
 For instance, when two user-provided objects are paired with a text prompt specifying a particular relationship, the generated output should not only preserve their identities but also accurately reflect the intended relationship, such as a ‘hug’.
 This introduces new challenges in what we refer to as relation-aware customized image generation, which focuses on preserving multiple identities while adhering to relationship prompts.
-An intuitive solution to this issue is to adapt existing tuning-based or training-based methods for customizing relationships (Huang et al., 2023; mengmeng Ge et al., 2024; Materzynska et al., 2023) among objects.
+An intuitive solution to this issue is to adapt existing relation-aware generation methods [11, 22, 21] to tuning-based or training-based customized methods for customizing objects while maintaining the relation.
 However, both approaches face challenges.
-Tuning-based methods struggle to preserve multiple identities, as they invert objects into specific tokens, while training-based schemes often fail to balance image and text prompts, frequently overlooking key textual elements and hindering the generation of relationships between objects.
+Tuning-based methods struggle to preserve multiple identities, as they invert objects into specific tokens.
+While training-based schemes often fail to balance image and text prompts, frequently overlooking key textual elements and hindering the generation of relationships between objects.
 As illustrated in Fig.
-1, previous methods fail to capture the actions described in the text prompt and lose identity preservation.
+1, previous methods fail to capture the relationships described in the text prompt and lose identity preservation.
 We attribute this failure to two key factors: a lack of relevant data and an ineffective model design.
-Unlike data augmentation techniques such as flipping or rotation, commonly used to create paired training data in object customization methods (Yang et al., 2023; Chen et al., 2024b), our approach requires a triplet of images: two image prompts and one target image.
-The image prompts should contain similar objects but exhibit distinct actions compared to the target image.
+Unlike data augmentation techniques such as flipping or rotation, commonly used to create paired training data in object customization methods [40, 5], our approach requires a triplet of images: two image prompts and one target image.
+The image prompts should contain similar objects but exhibit distinct poses compared to the target image.
 To collect these triplets, we propose a data engine to curate our fine-tuning set.
-We leverage an advanced text-to-image generation model (Betker et al., 2023) to generate triplets where the same object pair is shared across the images.
-Through text prompt guidance, the object in the image prompt provides strong identity information while maintaining distinct actions in the target image, enabling the model to focus on relation learning during fine-tuning.
-For the model design, we propose RelationBooth, which applies the Low-Rank Adaptation (LoRA) (Hu et al., 2022) strategy to the text cross-attention layers of existing diffusion models to process user-provided text prompts.
-In RelationBooth, two key modules are introduced during training to enhance the customization of relationships and identity preservation.
-First, we introduce a keypoint matching loss (KML) as additional supervision to explicitly encourage the model to adjust object poses, since relationships between objects are closely tied to their poses.
+We leverage an advanced text-to-image generation model [2] to generate triplets where the same object pair is shared across the images.
+Through text prompt guidance, the image prompt provides strong identity information.
+This enables the decoupling of the relationship in the target image, which enhances the relation learning.
+For the model design, we propose DreamRelation, which applies the Low-Rank Adaptation (LoRA) [9] strategy to the text cross-attention layers of existing diffusion models to process user-provided text prompts.
+In DreamRelation, two key modules are introduced during training to enhance the relation generation in customized generation.
+First, we introduce a keypoint matching loss (KML) as additional supervision to explicitly encourage the model to manipulate object poses, since relationships between objects are closely tied to their poses.
 Importantly, the KML operates on the latent representation rather than the original image space, aligning with the default diffusion loss.
-Second, we inject local tokens for multiple objects to improve the distinctiveness of highly overlapping objects.
-Specifically, we employ the self-distillation method from CLIPSelf (Wu et al., 2024b) to enhance region-language alignment in CLIP’s dense features.
-Through partitioning and pooling, the well-aligned local tokens help mitigate appearance confusion between objects.
-To more comprehensively evaluate relation-aware customized image generation, we constructed our RelationBench based on three established benchmarks (Ruiz et al., 2023; Kumari et al., 2023).
+Second, since relation-aware customization requires local features from image prompts, such as the “hands” features for generating “shaking hands”—which are not captured by CLIP’s coarse image-level features, we introduce dense features from CLIP.
+Through partitioning and pooling, we obtain local tokens that contain detail and local information.
+To further enhance the compatibility between dense features and image-level features, we employ a self-distillation method [39] to improve their alignment.
+To more comprehensively evaluate relation-aware customized image generation, we constructed our RelationBench based on three established benchmarks [31, 15].
 Our methods demonstrate strong performance compared to existing approaches, achieving significant improvements in visual quality and quantitative results.
 The contributions of this work are:
-  • We explore a novel task called relation-aware customized image generation, which aims to preserve multiple identities from image prompts while adhering to the relationships specified in text prompts.
-This task can enhance various user-driven applications by enabling more control and customization.
-  • We introduce a data engine that uses an advanced text-to-image generation model to generate triplet images where the same object pair is present with distinct actions.
-This well-curated dataset enables the model to focus on learning object relations during fine-tuning, leveraging the robust identity information provided by the image prompts.
-  • Our proposed RelationBooth method has two key modules, including keypoint matching loss and object-related local tokens, to enhance the customization of relationships and identity preservation.
+  • We explore a novel task called relation-aware customized generation, which aims to preserve multiple identities from image prompts while adhering to the relationships specified in text prompts.
+This task can enhance various user-driven applications by enabling more control.
+  • We introduce a data engine that uses the recent text-to-image generation model to generate triplet images where the same object is present with distinct poses.
+This well-curated dataset enables the model to focus on relation learning during fine-tuning, leveraging the identity information provided by the image prompts.
+  • Our proposed DreamRelation has two key modules, including keypoint matching loss and local token injection, to enhance the relation generation in customized generation.
 The extensive experiments on three benchmarks demonstrate the effectiveness of our method.
@@
 Diffusion-Based Text-to-Image Generation.
-This direction aims to produce high-quality images based on user-provided textual prompts.
-Diffusion-based generation models (Rombach et al., 2022; Saharia et al., 2022) encode these prompts through the text model (Radford et al., 2021; Chung et al., 2022) and inject the text embedding into the U-Net’s cross-attention layers.
-Some methods (Podell et al., 2023; Pernias et al., 2023; Ren et al., 2024; Wu et al., 2024a) upscale diffusion models and incorporate additional conditions as priors to generate high-resolution images.
-Meanwhile, several methods (Betker et al., 2023; Lian et al., 2023) introduce stronger text encoders or large language models (LLMs) to enhance the textual comprehension capabilities of diffusion models.
-Recent works (Peebles & Xie, 2023; Chen et al., 2024a) have replaced the U-Net denoiser with a Transformer-based denoiser.
-However, these models primarily focus on textual conditions and struggle to handle other forms of input, such as user-provided images.
-In contrast, our work facilitates customized generation while following predicate relation conditions given by text inputs.
+Diffusion-based text-to-image models [30, 32] generate high-quality images based on user-provided text prompts.
+These works encode the text prompt through the text encoder [28, 6] and inject the text embedding into the U-Net’s cross-attention layers.
+Some methods [27, 26, 29, 38] upscale diffusion models and incorporate additional conditions as priors to generate high-resolution images.
+Meanwhile, several methods [2, 17] introduce stronger text encoders or large language models (LLMs) to enhance the text comprehension capabilities of diffusion models.
+Recent works [25, 4] have replaced the U-Net denoiser with the Transformers [33].
+However, these models primarily focus on text conditions and struggle to handle other forms of guidance, such as user-provided images.
+In contrast, our work facilitates customized generation while following relation conditions given by text inputs.
 Diffusion-Based Customized Generation.
-Customized image generation aims to produce diverse images based on user-provided concepts.
-Tuning-based methods accomplish this by fine-tuning specific parameters of diffusion models, thereby incorporating new concepts into text-to-image diffusion models like Stable Diffusion.
-Some of these methods (Gal et al., 2023; Voynov et al., 2023; Liu et al., 2023) employ text embeddings to represent customized concepts, which are then injected into prompts during inference.
-On the other hand, Dreambooth (Ruiz et al., 2023) fine-tunes the entire U-Net and introduces a prior preservation loss to mitigate language drift.
-Additionally, several approaches (Kumari et al., 2023; Gu et al., 2024; Lin et al., 2024; Pang et al., 2024) combine text embeddings and diffusion models for image synthesis.
-They fine-tune both the text embedding and the U-Net, yielding impressive results.
-Considering the substantial cost of fine-tuning for commercial uses, training-based methods (Li et al., 2023; Wei et al., 2023; Chen et al., 2024b; Patel et al., 2024; Zhang et al., 2024) have been proposed.
-This approach primarily utilizes an ID encoder to extract the concept’s identity and inject it into the U-Net via cross-attention layers.
-For instance, MS-Diffusion (Wang et al., 2024) integrates grounding tokens with a feature resampler to maintain detail fidelity and employs layout guidance to explicitly place the concepts.
-Despite their efficiency during inference, training-based methods often struggle with identity preservation and textual condition following ability.
-A typical problem is the relation expressions in prompts are often overlooked.
-To address this issue, we present RealtionBooth, to tackle relation-aware customized generation.
+Customized image generation aims to produce new images based on user-provided objects.
+Tuning-based methods accomplish this by fine-tuning specific parameters of diffusion models to learn new objects.
+Some of these methods [7, 34, 20] employ text embeddings to represent the object identity.
+On the other hand, DreamBooth [31] fine-tunes the entire U-Net and introduces a prior preservation loss to mitigate language drift.
+Additionally, several approaches [15, 8, 18, 23] integrate text embeddings and U-Net parameters for image customization.
+These methods fine-tune both components during object learning, achieving impressive results.
+Considering the substantial cost of fine-tuning for commercial uses, training-based methods [16, 36, 5, 24, 43] have been proposed.
+This approach primarily utilizes an encoder to extract the object’s identity and inject it into the U-Net.
+For instance, MS-Diffusion [35] integrates grounding tokens with a Resampler to maintain detailed identity and employs layout guidance to explicitly locate the objects.
+Despite their efficiency during inference, training-based methods often face challenges in balancing identity preservation and text control.
+A common issue is that the relationships described in text prompts are frequently overlooked.
+To address this limitation, we introduce DreamRelation, a framework designed to bridge the customization and relation generation.
 Relation-Aware Text-to-Image Generation.
-Inspired by tuning-based customization methods (Gal et al., 2023), some recent works (Huang et al., 2023; Wei et al., 2024; mengmeng Ge et al., 2024) represent a “neglected word” by learnable parameters.
-These methods fine-tune part of the parameters on content co-existing images.
-For instance, Reversion (Huang et al., 2023) fine-tunes the text embedding on a set of co-existing related images and introduces a relation-steering contrastive loss.
+Inspired by tuning-based customization methods [7], some recent works [11, 37, 22] represent a “neglected word” by learnable parameters.
+These methods fine-tune part of the parameters on the content co-existing images.
+For instance, Reversion [11] fine-tunes the text embedding on a set of relation co-existing images and introduces a relation-steering contrastive loss.
 Reversion gains a better alignment between relational words and generated images by injecting new text embedding into prompts.
-While effective with prepositions and adjectives, Reversion struggles with predicate relation which involves significant overlaps, and cannot customize user-provided concepts.
-Moreover, this method still struggles to generate vivid relationships while maintaining the fidelity of multiple customized concepts, likely due to the inherent disregard for text embeddings in customization methods.
+While effective with prepositions and adjectives, Reversion struggles with relationships that involve significant overlaps, and cannot customize user-provided objects.
+Moreover, this method still struggles to generate vivid relationships while maintaining the fidelity of multiple customized objects, likely due to the inherent disregard for text embeddings in customization methods.
 Our work bridges the gap in the predicate relation-following capability of training-based customization methods.
-This enables efficient inference that naturally and accurately generates predicate relations from the textual prompt.
-3 Relation-aware customized obejct generation
-Figure 2: The overview of RelationBooth.
-RelationBooth encodes the image prompts by dual encoder architecture.
-After getting the U-Net output $\hat{\epsilon}$, we predicate $\hat{z}_{0}$ and calculate the keypoint matching loss.
+This enables efficient inference that naturally and accurately generates relationships in the text prompts.
+3 Bridging Customization and Relation Generation
+Figure 2: The overview of DreamRelation.
+DreamRelation utilizes the off-the-shelf identity extractor to decouple the relation and identity information in relation-specific images.
+After getting the U-Net output $\hat{\epsilon}$, we predict $\hat{z}_{0}$ and calculate the keypoint matching loss.
 The part in the dotted box is only for training.
 3.1 Problem definition
-Different from convention customization tasks, we explore a new setting that focuses on image generation by both image prompts $c_{i}\in\mathcal{R}^{N\times H\times W}$ and text prompts $c_{t}$.
-The $N$, $H$, and $W$ are the number, height, and width of image prompts.
-We call this setting RelationBooth due to the requirements that the generated image $\hat{x}$ should strictly preserve each identity and keep the predicate relation among those identities provided by $c_{i}$ and $c_{t}$.
+Different from conventional customization tasks, we explore a new setting that focuses on image generation by both image prompts $c_{i}\in\mathcal{R}^{N\times 3\times H\times W}$ and text prompts $c_{t}$.
+We call this setting relation-aware customized image generation due to the requirements that the generated image $\hat{x}$ should strictly preserve each identity and keep the relationship among those identities provided by $c_{i}$ and $c_{t}$.
 We define this task as:
 $\hat{x}=\Phi_{\theta}(c_{i},c_{t})$ (1)
-where $\Phi$ is the network parametered by $\theta$. For brevity clarification, we set $N=2$ due to the basic triplet element defined in the RelationBooth.
-3.2 Discussion and Motivation of RelationBooth
-Although previous works (Kumari et al., 2023; Gu et al., 2024; Zhang et al., 2024) have attempted to customize multiple objects under text control, we are the first to address the generation of relation-aware customized images.
-Notably, the state-of-the-art multi-object customization model, MS-Diffusion (Wang et al., 2024), cannot handle this task effectively.
-Specifically, given the image prompt $c_{i}$ and text prompt $c_{t}$.
-MS-Diffusion utilizes the CLIP (Radford et al., 2021) to extract the image and text tokens from $c_{i}$ and $c_{t}$ respectively.
-The image and text tokens are injected into the U-Net $\bm{\epsilon_{\theta}}$ through parallel cross-attention layers.
-The U-Net is optimized on the latent representation of image $z$:
-$\mathcal{L}_{MS}=\mathbb{E}_{z_{t},t,\epsilon,c}[\|\epsilon_{t}-\bm{\epsilon_{\theta}}(z_{t},t,c_{i},c_{t})\|_{2}^{2}]$ (2)
-where $\epsilon_{t}\sim\mathcal{N}(0,1)$, we omit the CLIP encoder for brevity.
-However, despite MS-Diffusion can customize objects under text control, it’s struggled for MS-Diffusion to generate relations accurately.
-Furthermore, MS-Diffusion might cause object confusion when heavy overlapping exists, leading to failure relation generation.
-We revise the relation generation task.
-Existing relation inversion methods (Huang et al., 2023; mengmeng Ge et al., 2024) aim to invert a relation to a text embedding $R^{*}$ in the pre-trained text-to-image diffusion models.
-Given a set of relation-specific images $\{x_{k}\}_{k=1}^{n}$, they employ denoising loss $\mathcal{L}_{RI}$ to fine-tune $R^{*}$ for alignment with specific relation:
+where $\Phi_{\theta}$ is the network parameterized by $\theta$. For brevity clarification, we set $N=2$ due to the basic triplet element defined in the DreamRelation.
+3.2 Discussion and Motivation
+While previous works [15, 8, 43] have explored the customization of multiple objects under text control, they lack the ability to effectively manage the relationships between these objects.
+Specifically, the state-of-the-art multi-object customization model [35] utilizes bounding boxes to enhance multi-object generation.
+Given the image prompt $c_{i}$ and the text prompt $c_{t}$, it leverages CLIP [28] and the Resampler [35] to extract text tokens $tok_{\text{text}}$ and image tokens $tok_{\text{image}}$, respectively.
+These tokens are then injected into the U-Net $\bm{\epsilon_{\theta}}$ through parallel cross-attention layers:
+$h=\text{Softmax}(\frac{QK_{i}}{\sqrt{d}}+\mathcal{M})V_{i}+\text{Softmax}(\frac{QK_{t}}{\sqrt{d}})V_{t}$ (2)
+where $Q$ represents the query, while $K_{i},V_{i}$ and $K_{t},V_{t}$ denote the keys and values obtained from $tok_{\text{image}}$ and $tok_{\text{text}}$, respectively.
+The mask $\mathcal{M}$, generated from the bounding boxes, assists the model in localizing objects during multi-object customization.
+However, it encounters difficulties when bounding boxes overlap, frequently resulting in object confusion, as shown in Fig.
+1.
+Consequently, relation-aware generation in customized models remains a largely unexplored area of research.
+On the other hand, we revisit the relation-aware generation task.
+Existing relation inversion methods [11, 22] focus on inverting a relationship into a text embedding $R^{*}$ within pre-trained text-to-image diffusion models.
+Given a set of images $\{x_{k}\}_{k=1}^{n}$ that share the same relationship, these methods use a denoising loss $\mathcal{L}_{RI}$ to fine-tune $R^{*}$, ensuring alignment with the specific relationship:
 $\mathcal{L}_{RI}=\mathbb{E}_{z_{t},t,\epsilon,c}[\|\epsilon_{t}-\bm{\epsilon_{\theta}}(z_{t},t,c_{t})\|_{2}^{2}]$ (3)
-Directly implementing relation inversion methods on MS-Diffusion leads to suboptimal results.
-We fine-tune MS-Diffusion on a set of relation-specific images, and experiment results show the model falls short in generating relation in Fig.
+Directly applying relation-aware generation methods to customized models leads to suboptimal performance, as illustrated in Fig.
 4.
-We think the image prompt $c_{i}$ that is not considered in relation inversion methods affects the generation process, leading to an overlook of $R^{*}$.
-In the following subsections, we will introduce the data collection and proposed method, considering lacking both of them in the proposed tasks.
+We attribute this to two main reasons.
+First, the balance between the control of $c_{i}$ and $c_{t}$ is not addressed in relation-aware generation tasks.
+During inference, the model tends to overlook the target relationship $R^{*}$ due to the influence of $c_{i}$, resulting in incorrect relationship generation.
+Second, even if the shared relationship exists across $\{x_{k}\}_{k=1}^{n}$, the disentangled object identity information complicates the learning process.
+Fine-tuning solely on $\{x_{k}\}_{k=1}^{n}$ makes it challenging for the model to effectively capture the relationship.
+As shown in the second row of Fig.
+4, where ReVersion [11] struggles to generate the ”hugging” relationship even in the absence of $c_{i}$.
+In the following subsections, we will introduce our data collection method and DreamRelation to bridge customization and relation generation.
 3.3 Data Collection
-Considering the scarcity of available datasets that can be used, we first design a data engine for collecting high-quality tuning data.
-We denote the relation-specific image as $x$.
-While $x$ contains relation information, it also includes the object’s identities.
-Directly fine-tuning on a set of $x$ leads to unsatisfied relation generation.
-To disentangle the relation and identity in $x$.
-The ideal tuning data should be in the form of $\mathcal{D}=(x,c_{i},c_{t})$.
-Where $c_{i}$ represents the image prompts that contain the object in $x$, $c_{t}$ represents the text prompt that guides the relation generation.
-Unfortunately, directly cropping the objects from $x$ to obtain $c_{i}$ results in a copy-and-paste effect, even when using data augmentation techniques like flipping and rotation.
+To address the balance between $c_{i}$ and $c_{t}$ while facilitating relation learning from $\{x_{k}\}_{k=1}^{n}$, we first design a data engine for collecting high-quality tuning data.
+The ideal tuning data should be in the form of $\mathcal{D}=(x_{k},c_{i},c_{t})$.
+The $c_{i}$ contain the object in $x_{k}$ while $c_{t}$ describes the relationship between objects in $x_{k}$.
+The identity information in $c_{i}$ helps decouple the relationship within $x_{k}$, enhancing relation learning while maintaining a balanced integration of $c_{i}$ and $c_{t}$.
+Unfortunately, directly cropping the objects from $x_{k}$ to obtain $c_{i}$ results in a copy-and-paste effect [5], even when using data augmentation techniques like flipping and rotation.
 Therefore, we introduce our data engine below.
-Data engine.
+Relation-aware data engine.
 Fig.
-3 illustrates our data generation process.
-Inspired by the capability of DALL-E 3 and its multi-turn dialog capability.
+3 illustrates the difference of our data engine.
+Both of $x_{k}$ and $c_{i}$ are generated by the recent text-to-image generation model, rather than through cropping and augmentation.
+Specifically, inspired by the performance of DALL-E 3 and its multi-turn dialog capability.
 We use DALL-E 3 to generate $\mathcal{D}=(x,c_{i},c_{t})$.
-Where $x$ and $c_{i}$ share the same object identity by leveraging the prompt “The photo of the same.” We observe that prompting in this manner allows DALL-E 3 to remember and preserve identity in common categories such as “tiger” and “brown bear” which are enough for relation learning.
-Next, we employ X-Pose (Yang et al., 2024), SAM (Kirillov et al., 2023), and LLaVA (Liu et al., 2024) to annotate $x$, $c_{i}$ with keypoints, masks, and captions.
-Figure 3: The overview of the data generation process. DALLE-3 can generate similar objects by using ’The photo of the same’ in the text prompt on common animal categories.
-Figure 4: Comparison of different fine-tuning data. Simply implementing relation inversion methods on MS-Diffusion leads to failure relation generation.
-3.4 RelationBooth
-3.4.1 RelationLoRA Fine-tuning
-To disentangle the relation and identity information in relation-specific images $x$, we utilize off-the-shelf identity inject methods from state-of-the-art customized model MS-Diffusion (Wang et al., 2024).
-Specifically, we utilize parallel cross-attention layers to process the $c_{i}$ and $c_{t}$, the identity information in $c_{i}$ facilitates the model to focus on relation information in $x$ and $c_{t}$ during fine-tuning.
-We formulate the parallel cross-attention layers:
-$h=\gamma\cdot\operatorname{Softmax}(\frac{QK_{i}}{\sqrt{d}})V_{i}+\operatorname{Softmax}(\frac{QK_{t}}{\sqrt{d}})V_{t}$ (4)
-where $Q=W_{q}h$, $K_{i}=W_{k_{i}}c_{i}$, $V_{i}=W_{v_{i}}c_{i}$, $K_{t}=W_{k_{t}}c_{t}$, $V_{t}=W_{v_{t}}c_{t}$, and $\gamma$ is a hyperparameter for scaling features from $c_{i}$.
+Where $x_{k}$ and $c_{i}$ share the same object identity by leveraging the prompt “The photo of the same.” We observe that prompting in this manner allows DALL-E 3 to remember and preserve identity in common categories such as ”tiger”, which is enough for relation learning.
+Next, we employ X-Pose [41], SAM [13], and LLaVA [19] to annotate $x$, $c_{i}$ with keypoints, masks, and captions for samples.
+Figure 3: The rigid object cropping data engine, which results in minimal changes to object pose, the cropped image prompts also contain relation information, leading to a copy-and-paste effect, where the text prompt is neglected.
+Our relation-aware data engine, on the other hand, focuses on relation learning by decoupling identity information from images.
+3.4 DreamRelation
+3.4.1 Relation Learning
+To balance $c_{i}$ and $c_{t}$ while enhancing relation learning, we decouple the identity information and relationships in $x_{k}$.
+Specifically, we leverage an off-the-shelf ID extractor from the state-of-the-art customized model [35].
+The identity information in $c_{i}$ enables the model to focus on the relationship information in $x_{k}$ and $c_{t}$ during relation learning.
+To achieve this, we employ parallel cross-attention layers to process $c_{i}$ and $c_{t}$, which can be formulated as:
+$h=\gamma\cdot\text{Softmax}(\frac{QK_{i}}{\sqrt{d}}+\mathcal{M})V_{i}+\text{Softmax}(\frac{QK_{t}}{\sqrt{d}})V_{t}$ (4)
+where $Q=W_{q}h$, $K_{i}=W_{k_{i}}c_{i}$, $V_{i}=W_{v_{i}}c_{i}$, $K_{t}=W_{k_{t}}c_{t}$, $V_{t}=W_{v_{t}}c_{t}$, and $\gamma$ is a hyperparameter for scaling control from $c_{i}$.
 For clarity, we omit the final linear layer $W_{out}$.
-The predicate relation information mainly depends on text prompts.
-Therefore, we inject LoRA layers in all of the $W_{q}$, $W_{k_{t}}$, $W_{v_{t}}$, and $W_{out}$ to encourage the model to pay more attention to the predicate relation in $c_{t}$.
+The relation information mainly depends on text prompts.
+Therefore, we inject LoRA layers in all of the $W_{q}$, $W_{k_{t}}$, $W_{v_{t}}$, and $W_{out}$ to encourage the model to pay more attention to the relationship in $c_{t}$.
 We freeze all other parameters during fine-tuning.
-The required additional storage space for LoRA weights is only 13MB.
-As shown in Fig.4, benefited from the disentangled identity information in $c_{i}$, the LoRA weights effectively capture the predicate relation in $x$ and $c_{t}$, accurately generating the corresponding images.
-Notably, after fine-tuning, our RelationLoRA can also be directly integrated into SDXL (Podell et al., 2023), effectively enabling the model to address the relation inversion task.
-The extensive results are presented in the appendix (Fig.
-15).
+As shown in Fig.
+4, benefiting from the decoupled identity information in $c_{i}$, our method effectively captures the relationship in $x_{k}$ and $c_{t}$, accurately generating the corresponding images.
+Notably, after fine-tuning, our method can also be directly integrated into SDXL [27], effectively enabling the model to address the relation-aware generation task.
+The extensive results are presented in the Supplementary Material (SM).
+Methods Single-Object Cus Multi-Object Cus Non-Confusing Relation-Aware
+Textual Inversion [7]
+DreamBooth [31]
+Custom Diffusion [15]
+Mix-of-Show [8]
+CLIF [18]
+MultiBooth [45]
+BLIP-Diffusion [16]
+ELITE [36]
+AnyDoor [5]
+SSR-Encoder [43]
+$\lambda$-eclipse [24]
+MS-Diffusion [35]
+ReVersion [11]
+ADI [10]
+DreamRelation
+Table 1: Setting Comparison for Different Models. We include several representative methods here. Our proposed DreamRelation can meet the demand of the relation-aware customization task.
+Figure 4: Comparing our method with ReVersion across different base models, our approach demonstrates superior performance in the relation-aware generation task.
 3.4.2 Keypoint Matching Loss
-Most relations have specific requirements for the pose, such as “hugging” requires the arm to crossed over the other object’s body, and “riding a bicycle” requires the feet of the object to be on the pedal plate.
-Therefore, it’s important to accurately manipulate the pose of objects in the right status.
-Intuitively, we propose to introduce keypoint features into the fine-tuning procedure to explicitly guide the pose of the objects for better relation generation.
-Specifically, considering our task involves common objects instead of humans only, We employ X-Pose (Yang et al., 2024) as keypoint detector because it can detect any keypoints in complex real-world scenarios.
-We detect 17 keypoints for each object in $x$ and $c_{i}$, and denote the keypoint coordinates as $c^{x}_{kp},c^{c_{i}}_{kp}\in\mathcal{R}^{17\times 2}$ respectively.
-To encourage the model to generate accurate pose in $\mathcal{D}(\hat{z}_{0})$, where $\mathcal{D}$ is the VAE decoder.
+Most relationships have specific requirements for the pose, such as ”hugging” requiring the arms to cross over the object’s body, and ”riding a bicycle” requiring the feet of the object to be on the pedal plate.
+Therefore, it’s important to manipulate the pose of objects in the right status accurately.
+Intuitively, we introduce explicit supervision in the diffusion latent space during the tuning process, named Keypoint Matching Loss (KML).
+Specifically, considering our task only involves common objects instead of humans, we employ X-Pose [41] as a keypoint detector because it can detect any keypoints in complex real-world scenarios.
+We detect 17 keypoints for each object in $x_{k}$ and $c_{i}$, and denote the keypoint coordinates as $c^{x_{k}}_{kp},c^{c_{i}}_{kp}\in\mathcal{R}^{17\times 2}$ respectively.
+To encourage the model to generate accurate poses in $\mathcal{D}(\hat{z}_{0})$, where $\mathcal{D}$ is the VAE decoder.
 We use the U-Net’s output $\hat{\epsilon}$ to predict $z_{0}$ during fine-tuning:
 $\hat{z}_{0}=\frac{z_{t}-\sqrt{1-\bar{\alpha}_{t}}\hat{\epsilon}}{\sqrt{\bar{\alpha}_{t}}}$ (5)
-We use the VAE encoder $\mathcal{E}$ to obtain latent representation of image prompts $\mathcal{E}(c_{i})$.
+We use the VAE encoder $\mathcal{E}$ to obtain the latent representation of image prompts $\mathcal{E}(c_{i})$.
 Then we calculate the MSE loss on the corresponding keypoint locations between $\mathcal{E}(c_{i})$ and $\hat{z}_{0}$:
-$\mathcal{L}_{KML}$ $=\frac{1}{n_{kp}}\mathbb{E}_{\bm{z_{t}},c_{i}}\left\|\mathcal{E}(c_{i})[c^{c_{i}}_{kp}]-\hat{z_{0}}[c^{z_{0}}_{kp}]\right\|^{2}_{2}$ (6)
+$\mathcal{L}_{KML}$ $=\frac{1}{n_{kp}}\mathbb{E}_{\bm{z_{t}},c_{i}}\left\|\mathcal{E}(c_{i})[c^{c_{i}}_{kp}]-\hat{z}_{0}[c^{x_{k}}_{kp}]\right\|^{2}_{2}$ (6)
 $\mathcal{L}$ $=\mathcal{L}_{denoise}+\lambda\cdot\mathcal{L}_{KML}$ (7)
-where $n_{kp}$ denote the number of keypoint, $c^{c_{i}}_{kp}$ and $c^{z_{0}}_{kp}$ are keypoint’s coordinates in $c_{i}$ and $z_{0}$ respectively, $\lambda$ controls for the relative weight of KML.
+where $n_{kp}$ denote the number of keypoint, $c^{c_{i}}_{kp}$ and $c^{x_{k}}_{kp}$ are keypoint’s coordinates in $c_{i}$ and $x_{k}$ respectively.
+We scale the coordinates to fit the size of $\mathcal{E}(c_{i}),\,\hat{z}_{0}$ $\lambda$ controls for the relative weight of KML.
 The part inside the dotted box in Fig.
 2 illustrates the model fine-tuning with the KML.
-We find KML is effective in encouraging the model to manipulate the pose for relation generation.
+We find KML is effective in encouraging the model to manipulate the pose for relation-aware generation.
+Figure 5: Additional results demonstrate the effectiveness of relation-aware generation. DreamRelation adapts the same pair of objects to different relationships in a natural and accurate manner.
 3.4.3 Local Token Injection
-When generating relations, overlapping objects often occur.
-It is challenging to distinguish multiple objects solely using CLIP’s image-level embedding of the full image, which can lead to confusion between customized objects during relation generation.
-To address this limitation, we introduce a local region representation of images.
-Considering the domain shift from the full image to local image regions, we employ CLIPSelf’s self-distillation methods (Wu et al., 2024b) to enhance the region-language alignment of local region representation.
-Specifically, in line with CLIPSelf, we extract dense features (Zhou et al., 2022) from a student CLIP model and derive local tokens through partitioning and pooling (He et al., 2017).
-We partition the full image into patches and feed them into a frozen CLIP model to obtain teacher tokens.
-Our approach minimizes the cosine similarity between local tokens and teacher tokens to improve the region-language alignment of local tokens.
-To formulate this process, we denote the hidden states of the second-to-last layer of CLIP as $h^{\prime}_{clip}$.
-We modify the last layer of CLIP $h_{clip}=\operatorname{ModifiedAttention}(h_{clip})$ to obtain the dense features:
-$h_{tmp}=\operatorname{Proj}_{v}(\operatorname{norm}(h^{\prime}_{clip})),h_{tmp}=h^{\prime}_{clip}+\operatorname{Proj}_{out}(h_{tmp}),h_{tmp}=h_{tmp}+\operatorname{FFN}(h_{tmp})$ (8)
-where $\operatorname{Proj}$ means the linear projection.
-We discard the class embedding $h_{clip}[0]$ and reshape the image embeddings $z^{\prime}[1:h\times w]$ into an $h\times w$ feature map.
-We partition the feature map into patches, from which we can get local tokens $tok_{l}$ by pooling.
-During inference, we simply concatenate the aligned local tokens with the image-level tokens and send them to the ID extractor which is a transformer-based architecture shown in Fig.
-2 b).
-The query $q$ is obtained by concatenating class text embedding with the bounding box.
-$q=q+\operatorname{GatedSelfAttn}([q,tok_{i},tok_{l}])$ (9)
-Experimental results show that local token injection enhances identity preservation in evaluation metrics and visual effects as shown in Fig. 9.
+When generating relationships, local and detailed features are essential—for instance, the feature of arms to accurately construct relationships like “carrying”.
+However, the coarse image-level features extracted by the CLIP Image Encoder fail to provide such fine-grained details, leading to confusion between objects in Fig.
+7.
+To address this limitation, we introduce dense features from the CLIP Image Encoder [44].
+Specifically, we modify the last layer of the CLIP Image Encoder to obtain dense features: $h_{\text{dense}}=\text{ModifiedAttention}(h^{\prime}_{\text{clip}})$, where $h^{\prime}_{\text{clip}}$ represents the hidden states of the second-to-last layer.
+$h_{\text{tmp}}$ $=\text{Proj}_{v}(\text{norm}(h^{\prime}_{\text{clip}})),$ (8)
+$h_{\text{tmp}}$ $=h^{\prime}_{\text{clip}}+\text{Proj}_{out}(h_{\text{tmp}}),$ (9)
+$h_{\text{dense}}$ $=h_{\text{tmp}}+\text{FFN}(h_{\text{tmp}})$ (10)
+To further enhance the compatibility between dense features and image-level features, we employ self-distillation on the CLIP Image Encoder [39] to derive our Local Image Encoder.
+Specifically, we align local regions of the dense features with their corresponding image-level features using cosine similarity.
+During inference, we extract local tokens $tok_{\text{local}}$ from the dense feature $h_{\text{dense}}$ through partitioning and pooling.
+These local tokens are then concatenated with the image-level tokens $tok_{\text{image}}$ and passed to the ID extractor.
+The ID extractor, which is a transformer-based architecture as illustrated in Fig.
+2 (b), can be written as:
+$q=q+\text{Attention}(\text{concat}[q,\,tok_{\text{image}},\,tok_{\text{local}}])$ (11)
+Regarding the compatibility between CLIP dense features and the original image-level features, experimental results demonstrate that our Local Image Encoder outperforms other dense representations [3].
+For a detailed discussion, we refer readers to the Supplementary Material.
+Additionally, the injection of local tokens enhances identity preservation, as evidenced by improvements in evaluation metrics and visual results, shown in Fig.
+7 and Tab.
+3.
 4 Experiments
-Figure 5: Single-object comparison with our base model MS-Diffusion: The results demonstrate that our method generates more accurate and natural relation-aware images.
-Table 1: Quantitative comparison on RelationBench, Bold and underline represent the highest and second-highest metrics.
+Figure 6: We compare multi-object performance with training-based and tuning-based methods.
+Additionally, we integrate ReVersion [11] into MS-Diffusion, where the learnable text token is injected into the text prompt during MS-Diffusion’s inference, resulting in ReVersion+MS.
+Our method shows a clear advantage in relation-aware generation and avoiding object confusion in overlapping scenarios.
 Method Single-object Multi-object
 CLIP-T CLIP-R CLIP-I DINO CLIP-T CLIP-R CLIP-I DINO
-DreamBooth (SDXL) 27.8 18.2 74.2 62.6 24.3 16.2 67.8 57.2
-Custom Diffusion 26.5 15.2 73.2 58.8 20.1 15.4 64.7 55.3
-Cones-V2 24.4 13.5 72.1 57.2 21.3 15.2 64.3 54.2
-ELITE 25.7 14.9 75.4 61.5
-AnyDoor 24.5 14.7 77.4 62.2 21.6 14.9 69.7 59.8
-BLIP-Diffusion 26.2 15.7 77.4 57.7
-SSR-Encoder 25.5 15.9 80.4 59.4 24.2 14.6 72.1 56.2
-MS-Diffusion 26.5 18.8 78.7 64.5 26.9 18.9 73.8 58.8
+DreamBooth (SDXL) [31] 27.8 18.2 74.2 62.6 24.3 16.2 67.8 57.2
+Custom Diffusion [15] 26.5 15.2 73.2 58.8 20.1 15.4 64.7 55.3
+Cones-V2 [20] 24.4 13.5 72.1 57.2 21.3 15.2 64.3 54.2
+ELITE [36] 25.7 14.9 75.4 61.5
+AnyDoor [5] 24.5 14.7 77.4 62.2 21.6 14.9 69.7 59.8
+BLIP-Diffusion [16] 26.2 15.7 77.4 57.7
+SSR-Encoder [43] 25.5 15.9 80.4 59.4 24.2 14.6 72.1 56.2
+MS-Diffusion [35] 26.5 18.8 78.7 64.5 26.9 18.9 73.8 58.8
+Reversion+MS [11] 27.8 19.3 77.8 63.1 27.2 19.2 74.2 59.7
 Ours 30.6 21.4 77.9 63.4 28.9 20.4 75.4 62.1
-Figure 6: Multi-object comparison with training-based and tuning-based methods: Compared to our base model, our method shows a clear advantage in relation-aware generation and avoiding object confusion in overlapping scenarios.
-Table 2: DreamBench.
+Table 2: Quantitative comparison on RelationBench, Bold and underline represent the highest and second-highest metrics.
+Figure 7: Ablation study of keypoint matching loss (KML) and local tokens: The images are generated using the same random seed. We use green and red boxes to highlight the main differences.
+Figure 8: DreamRelation for three objects. Compared to the baseline model, DreamRelation shows a clear advantage.
+Method Multi-object
+CLIP-T CLIP-R CLIP-I DINO
+w/o Relation-aware Data 27.3 19.4 75.3 59.8
+w/o Local Token Injection 28.5 19.5 75.1 59.9
+w/o Keypoint Matching Loss 27.4 19.2 75.2 61.2
+Full Model 28.9 20.4 75.4 62.1
+Table 3: Ablation study on our proposed components
+Implementation Details.
+We generate a tuning set for relation learning with 4-6 samples per relationship.
+We incorporate LoRA layers into all text cross-attention layers of the U-Net.
+We set the LoRA rank to $r=4$, the parallel cross-attention scaling factor to $\gamma=0.6$, and the keypoint matching loss weight to $\lambda=1e-3$.
+The model is fine-tuned for 500 steps, using 2 A100 GPUs, with a total batch size of 8, completing the process in 10 minutes.
+We use the Adam optimizer [12] with a learning rate of $41e-$ 4 and no weight decay, resulting in a total of 3.1M trainable parameters.
+Note that our learnable parameters are only in the text cross-attention layers, making it compatible with any SDXL-based models.
+Considering the strong identity preservation capabilities of MS-Diffusion, we implement our DreamRelation on MS-Diffusion.
+During fine-tuning and inference, we concatenate local tokens with image tokens.
+We put more self-distillation details of the Local Image Encoder in the Supplementary Material.
+Evaluation.
+To evaluate relation-aware customized image generation, we propose RelationBench, consisting of 44 objects from DreamBench [31] and CustomConcept101 [15], along with 25 relationships.
+The object categories include pets, plushies, toys, people, and cartoons.
+Using GPT-4, we generate 100 cases for single-object and multi-object evaluations.
+Following previous works [31, 35], we evaluate our method on three metrics: (1) Identity Preservation, which assesses the alignment between the generated images and the image prompts, using the CLIP image score and DINO score that calculate the cosine similarity between the class embeddings of images, referred to as CLIP-I and DINO, respectively; (2) Text Alignment, which evaluates how well the generated images align with the text prompts, using the CLIP image-text score, denoted as CLIP-T; and (3) Relation Alignment.
+We observed that nouns in the prompts can inflate CLIP-T scores, making them less accurate for evaluating relation generation.
+To address this, we extract the relation word from the text prompt using spaCy and calculate the CLIP image-text score, denoted as CLIP-R.
+4.1 Main Results
+We evaluate our method through both quantitative and qualitative results in single- and multi-object cases.
+In single-object cases, we primarily focus on the alignment of the object’s pose with the relationship in the text prompt and the model’s ability to preserve identity.
+In multi-object cases, in addition to relation generation and identity preservation, we also need to ensure no confusion between the multiple objects.
+Quantitative results.
+We compare our method with baseline models across three benchmarks: RelationBench, DreamBench, and multi-object cases in CustomConcept101.
+As shown in Tab.
+2, our method demonstrates a clear advantage on RelationBench for single-object cases, particularly in the CLIP-T and CLIP-R metrics.
+However, its performance is slightly lower in the CLIP-I metric, likely due to significant pose variations in the objects, which may decrease the CLIP-I score.
+For multi-object generation, our method outperforms other approaches across all evaluation metrics, which we attribute to its ability to effectively avoid object confusion and significantly enhance CLIP-I performance.
+Furthermore, as shown in Tab.
+4, we evaluate single-object generation performance on DreamBench.
+Our method outperforms others in the CLIP-T metric while delivering competitive results in CLIP-I and DINO.
+We conduct experiments on multi-object cases from CustomConcept101, which we denote as M-CustomConcept101.
+As shown in Fig.
+5, our model achieves the highest performance in the DINO score and ranks second in both the CLIP-T and CLIP-I metrics.
+Qualitative Comparison.
+We conduct qualitative comparison experiments on RelationBench to evaluate relation-aware customized image generation.
+As shown in Fig.
+6, our method outperforms others in generating multi-object relationships, effectively avoiding object confusion while accurately capturing the intended relationships.
+Additional results can be found in the SM.
+4.2 Ablation study and Analysis
+Relation-Aware Data Engine.
+Instead of fine-tuning on our curated $(x_{k},c_{i},c_{t})$ pairs.
+We train a text embedding $R^{*}$ using a set of relation-specific data $(x_{k},c^{\prime}_{i},c_{t})$, where $c^{\prime}_{i}$ is cropped from $x_{k}$.
+The well-trained text embedding $R^{*}$ is injected into the text embedding during inference.
+The quantitative results in Tab.
+3 show that our relation-aware data engine outperforms in both CLIP-T and CLIP-R scores.
+Keypoint Matching Loss (KML).
+We omit KML during fine-tuning.
+As shown in Fig.
+7 a), including the KML leads to more accurate and natural relation-aware customized images using the same random seed.
+Quantitative results, presented in Tab.
+3, show a decline across all evaluation metrics when the KML is removed.
+Additionally, we experimented with different values for $\lambda$ and found that the best performance occurs at $\lambda=1e-3$.
+Local Token Injection.
+For local token injection, the visual comparison in Fig.
+7 b) shows severe object confusion when local tokens are omitted.
+The quantitative results in Tab.
+3 demonstrate that local tokens improve relation-aware customization.
+We conducted an ablation study on the number of local tokens, as presented in Tab.
+6, and determined the optimal number to be 16.
+Additional ablation studies are shown in the SM.
+Combining these methods ensures that DreamRelation achieves high-quality relation-aware customized generation.
 Method Single-object
 CLIP-T CLIP-I DINO
-DreamBooth (SDXL) 31.2 81.5 69.2
-Custom Diffusion 28.4 77.2 66.8
-Cones-V2 31.0 76.5 67.2
-ELITE 29.8 77.4 62.5
-AnyDoor 25.5 82.1 67.8
-SSR-Encoder 30.8 82.1 61.2
-MS-Diffusion 31.5 79.3 68.2
-Ours 31.6 78.9 67.4
-Table 3: M-CustomConcept101.
+DreamBooth (SDXL) [31] 31.2 81.5 69.2
+Custom Diffusion [15] 28.4 77.2 66.8
+Cones-V2 [20] 31.0 76.5 67.2
+ELITE [36] 29.8 77.4 62.5
+AnyDoor [5] 25.5 82.1 67.8
+SSR-Encoder [43] 30.8 82.1 61.2
+MS-Diffusion [35] 31.5 79.3 68.2
+Ours 31.6 81.7 68.4
+Table 4: Evaluation Results on DreamBench.
 Method Multi-object
 CLIP-T CLIP-I DINO
-DreamBooth (SDXL) 29.5 67.4 49.2
-Custom Diffusion 28.1 66.2 48.8
-Cones-V2 29.2 66.5 47.2
-AnyDoor 20.2 72.1 51.2
-SSR-Encoder 30.6 71.1 52.2
-$\lambda$-eclipse 29.2 68.2 48.2
-MS-Diffusion 28.0 70.2 51.2
-Ours 29.4 71.4 52.3
-Figure 7: Additional results of relation-aware generation.
+DreamBooth (SDXL) [31] 29.5 67.4 49.2
+Custom Diffusion [15] 28.1 66.2 48.8
+Cones-V2 [20] 29.2 66.5 47.2
+AnyDoor [5] 20.2 72.1 51.2
+SSR-Encoder [43] 30.6 71.1 52.2
+$\lambda$-eclipse [24] 29.2 68.2 48.2
+MS-Diffusion [35] 28.0 70.2 51.2
+Ours 29.7 71.4 52.3
+Table 5: Evaluation Results on M-CustomConcept101.
+Num Local Tokens Multi-object
+CLIP-T CLIP-R CLIP-I DINO
+N=2$\times$2 28.3 19.4 75.3 60.1
+N=4$\times$4 28.9 20.4 75.4 62.1
+N=8$\times$8 27.9 18.1 74.1 59.4
+Table 6: Ablation study on the number of local tokens.
+Lambda Multi-object
+CLIP-T CLIP-R CLIP-I DINO
+$\lambda$=1e-2 27.5 19.3 72.6 59.6
+$\lambda$=1e-3 28.9 20.4 75.4 62.1
+$\lambda$=1e-4 26.8 18.1 73.9 60.4
+Table 7: Ablation study on $\lambda$, controlling relative weight of KML.
+5 Conclusion
+In this work, we introduce a new challenging task: relation-aware customized image generation.
+This task aims to generate objects that adhere to the relationship specified in the text prompt and preserve the identities of the user-provided images.
+To support this, we propose a data engine for generating high-quality fine-tuning data.
+Our method, DreamRelation, incorporates Keypoint Matching Loss and Local Token Injection, effectively capturing relation information and generating natural relations between customized objects.
+We also present a fair comparison with other methods on a new benchmark, RelationBench.
+Extensive experiments demonstrate the effectiveness of our method, highlighting its potential for applications in interactive scenario generation, relation detection, and more.
+Our research can inspire the direction of relation-aware generation in the community.
+Acknowledgement. This work is supported by the National Key Research and Development Program of China (No. 2023YFC3807600).
+Supplementary Material
+Overview. The supplementary includes these sections:
+  • Sec. 6. Introduction video.
+  • Sec. 7. More experimental results.
+  • Sec. 8. Implementation details of the experiments.
+  • Sec. 9. Failure cases and limitations of our model.
+  • Sec. 10. Details of the RelationBench.
+  • Sec. 11. Incorporate with CogVideoX-I2V.
+  • Sec. 12. Social impacts.
+6 Introduction Video
+We provide a video introduction of our work. Please refer to “introduction_video.mp4” in the supplementary file.
+7 Additional Experimental Results
 Figure 9: Additional results of relation-aware generation across a wide range of objects.
-Figure 9: Ablation study of keypoint matching loss (KML) and local tokens: The images are generated using the same random seed.
-We use green and red boxes to highlight the main differences.
-The areas within the green boxes show the improvements with our proposed component.
-Table 4: Ablation study on our proposed components
-Method Single-object Multi-object
-CLIP-T CLIP-R CLIP-I DINO CLIP-T CLIP-R CLIP-I DINO
-w/o Relation-aware Data 28.4 16.9 78.2 64.4 26.2 16.8 75.3 59.7
-w/o Local Token Injection 29.2 21.0 76.8 62.5 28.5 19.5 75.1 59.3
-w/o Keypoint Matching Loss 28.3 17.9 77.8 62.1 27.4 18.2 75.2 61.2
-Full Model 30.6 21.4 77.9 63.4 28.9 20.4 75.4 62.1
+Keypoint Matching Loss.
+We use X-Pose [41] as our keypoint detector due to its open-vocabulary detection capabilities that are compatible with a wide range of objects instead of humans only.
+The keypoint matching loss (KML) facilitates relation generation by explicitly guiding the model’s pose manipulation, resulting in more accurate pose generation.
+As shown in Fig.
+11, the cat’s arm crosses the dog’s body, accurately depicting the ”hug” relation.
+The visualization in Fig.
+10 shows the cosine similarity between the latent representation of the image prompt and model prediction.
+Implementation with KML shows a better alignment of specific parts during training and inference.
+Figure 10: KML enhances the alignment of specific parts of the object in image prompt and model prediction.
+Figure 11: Additional ablation studies on KML and Local Tokens.
+Local Tokens Injection.
+To understand why local features enhance relation-aware generation, we adopt Principal Component Analysis (PCA) to project the dense feature more compactly.
+As shown in Fig.
+12, dense features provide more fine-grained information than CLIP image tokens and provide detailed information about each object to construct the interaction in relation.
+Aiding in distinguishing between different objects during the generation process and helping to avoid object confusion, especially in cases of heavy overlap.
+Moreover, it can facilitate object appearance alignment.
+For the injection method, we note that simply concatenating local tokens with image-level tokens produced the best performance as shown in Fig.
+9.
+Figure 12: Visualization of Dense feature and Image tokens by Principal Component Analysis (PCA).
+Figure 13: Multi-object relation-aware image customization results of pet, toy, plushie, and person.
 Table 8: Ablation study on Local Image Encoder’s architecture.
@@
 CLIP-ViT-bigG 28.9 20.4 75.4 62.1
-Table 6: Ablation study on $\lambda$, controlling relative weight of KML.
-Lambda Multi-object
-CLIP-T CLIP-R CLIP-I DINO
-$\lambda$=1e-2 27.5 19.3 72.6 59.6
-$\lambda$=1e-3 28.9 20.4 75.4 62.1
-$\lambda$=1e-4 26.8 18.1 73.9 60.4
-Table 7: Ablation study on the number of local tokens (determined by partition size).
-Num Local Tokens Multi-object
-CLIP-T CLIP-R CLIP-I DINO
-N=2$\times$2 28.3 19.4 75.3 60.1
-N=4$\times$4 28.9 20.4 75.4 62.1
-N=8$\times$8 27.9 18.1 74.1 59.4
-Table 8: Ablation study on local token injection methods to ID extractor.
+Table 9: Ablation study on local token injection methods.
 Injection method Multi-object
@@
 Concatenate 28.9 20.4 75.4 62.1
-Implementation Details.
-For relation-aware fine-tuning, we generate a tuning set with 4-6 samples per relation.
-We incorporate LoRA layers into all text cross-attention layers of the U-Net.
-We set the LoRA rank to $r=4$, the parallel cross-attention scaling factor to $\gamma=0.6$, and the keypoint matching loss weight to $\lambda=1e-3$.
-The model is fine-tuned for 500 steps, using 2 A100 GPUs, with a total batch size of 8, completing the process in 10 minutes.
-We use the Adam optimizer with a learning rate of $41e-$ 4 and no weight decay, resulting in a total of 3.1M trainable parameters.
-Note that our RelationLoRA only modifies the text cross-attention layers in SDXL, making it compatible with any SDXL-based models.
-Considering the strong identity preservation capabilities of MS-Diffusion, we implement our RelationBooth on MS-Diffusion.
-During fine-tuning and inference, we concatenate local tokens with image tokens.
-We put more self-distillation details of the Local Image Encoder in appendix (6.1).
-Additionally, as shown in appendix (Fig.
-15), our RelationLoRA is compatible with SDXL for addressing the Relation Inversion task (Huang et al., 2023).
-Evaluation.
-To evaluate relation-aware customized image generation, we propose RelationBench, consisting of 44 objects from DreamBench (Ruiz et al., 2023) and CustomConcept101 (Kumari et al., 2023), along with 25 predicate relations.
-The object categories include pets, plushies, toys, people, and cartoons.
-Using GPT-4, we generate 100 cases for single- and multi-object evaluations.
-Following previous works (Ruiz et al., 2023; Wang et al., 2024), we evaluate our method on three metrics: (1) Identity Preservation, which assesses the similarity between the generated images and the image prompts, using the CLIP image score and DINO score that calculate the cosine similarity between the class embeddings of images, referred to as CLIP-I and DINO, respectively; (2) Text Alignment, which evaluates how well the generated images align with the text prompts, using the CLIP image-text score, denoted as CLIP-T; and (3) Relation Alignment.
-We observed that nouns in the prompts can inflate CLIP-T scores, making them less accurate for evaluating relation generation.
-To address this, we extract relation predicate from the prompts using spaCy and calculate the CLIP image-text score using only the predicate, denoted as CLIP-R.
-4.1 Main Results
-We evaluate our method through both quantitative and qualitative results in single- and multi-object cases.
-In single-object cases, we primarily focus on the alignment of the object’s pose with the predicate relation in the text prompt, as well as the model’s ability to preserve identity.
-In multi-object cases, in addition to relation generation and identity preservation, we also need to ensure there is no confusion between the multiple objects.
-Quantitative results.
-We quantitatively compare our method with baseline models across three benchmarks: RelationBench, DreamBench, and multi-object cases in CustomConcept101.
+Relation-Aware Customization Data Engine.
+Our relation-aware data are generated by DALLE-3 [2], which can effectively preserve the identity of some categories through its multi-turn dialogue capability.
+Specifically, we preserve the object’s identity by adding ”the photo of the same” in the text prompt.
+Although our relation-aware data only contained animal categories, it can effectively learn the relation information in images and generalize to a wide range of objects as shown in Fig.
+9.
+Figure 14: Our fine-tuning dataset as an example.
+Relation-Aware Customized Image Generation.
+We show more qualitative results in Fig.
+16, Fig.
+21, and Fig.
+9.
+We compare our methods with both training-based and tuning-based methods.
+Our method shows a clear advantage in pose manipulation to fit the relation.
+We conduct additional results in Fig.
+15 to show DreamRelation’s differences with our base model-MS-Diffusion.
+Figure 15: Single-object comparison with our base model MS-Diffusion: The results demonstrate that our method generates more accurate and natural relation-aware images.
+Figure 16: Single-object comparison. TI and DB indicate Textual Inversion and DreamBooth, respectively. Our methods achieve the best balance between relation generation and identity preservation.
+Relation Inversion Task.
+As shown in Fig.
+17, our DreamRelation enhances SDXL’s ability to generate images that accurately adhere to specific relations.
+Compared to ReVersion [11], our method generates more precise relations, without object confusion or missing, demonstrating the strong performance of DreamRelation in the Relation Generation task.
+Figure 17: Our DreamRelation is compatible with SDXL to address Relation Inversion task.
+Evaluation on DreamBench and CustomConcept101.
 As shown in Tab.
-1, our method demonstrates a clear advantage on RelationBench for single-object cases, particularly in the CLIP-T and CLIP-R metrics.
-However, it performs slightly lower in the CLIP-I metric, likely due to significant pose variations in the objects, which can negatively impact the CLIP-I score.
-For multi-object generation, our method surpasses other approaches across all evaluation metrics, which we attribute to effectively preventing object confusion, leading to a significant improvement in CLIP-I performance.
-Furthermore, as shown in Tab.
-3, we evaluate single-object generation performance on DreamBench.
+4 and Tab.
+5, we evaluate single-object generation performance on DreamBench.
 Our method outperforms others in the CLIP-T metric while delivering competitive results in CLIP-I and DINO.
@@
 As shown in Fig.
-3, our model achieves the highest performance in the DINO score and ranks second in both the CLIP-T and CLIP-I metrics.
-Qualitative Comparison.
-We conduct qualitative comparison experiments for relation-aware customized image generation on RelationBench.
-For single-object comparisons, as shown in Fig.
-5, our generated images are more accurate than those produced by our base model, MS-Diffusion, and naturally reflect the predicate relations in the text prompts.
-Additional single-object comparison results are provided in the appendix (Fig.
-12).
-In multi-object comparisons, Fig.
-6 demonstrates that our method excels in generating multi-object relations, successfully avoiding object confusion while accurately representing the intended relations.
-4.2 Ablation study and Analysis
-RelationLoRA Fine-tuning.
-Instead of injecting LoRA layers and fine-tuning on our curated $(x,c_{i},c_{t})$, we fine-tune the text embedding $R^{*}$ using a set of relation-specific data $(x,c_{t})$, where blank images are used as image prompts $c_{i}$.
-The visual comparison is presented in Fig.
-4, and the quantitative results in Tab.
-4 indicate that this method underperforms in both CLIP-T and CLIP-R scores.
-This may be due to the entanglement of identity and relation information in $x$, as well as the gap between the blank and real images, making it difficult to accurately generate the relation.
-Keypoint Matching Loss.
-For the keypoint matching loss (KML), we omit it during fine-tuning.
-As shown in Fig.
-9 a), including the KML leads to more accurate and natural relation-aware images when using the same random seed.
-Quantitative results, presented in Tab.
-4, show a decline across all evaluation metrics when the KML is removed.
-Additionally, we experimented with different values for $\lambda$ and found that the best performance occurs at $\lambda=1e-3$.
-Local Token Injection.
-For local token injection, the visual comparison in Fig.
-9 b) shows severe object confusion when local tokens are omitted.
-The quantitative results in Tab.
-8 demonstrate that local tokens improve relation-aware customization.
-We conduct further ablation as shown in Tab.
-8, Tab.
-6, and Tab.
-8.
-We experimented with varying numbers of local tokens injected into the ID extractor and found that 16 tokens yielded the best results across all evaluation metrics.
-Therefore, we use 16 local tokens in all experiments.
-Additionally, we swapped the Local Image Encoder and found that CLIP-ViT-bigG was the most compatible model.
-For the injection method, we note that simply concatenating local tokens with image tokens produced the best outcomes.
-Combining these methods ensures that our method maintains high quality relation-aware customized generation.
-5 Conclusion
-In this work, we introduce a challenging task: relation-aware customized image generation, which aims to generate objects that adhere to the relations specified in the text prompt while preserving the identities of user-provided images.
-To support this, we propose a data engine for generating high-quality fine-tuning data.
-Our method, RelationBooth, incorporating Keypoint Matching Loss and Local Token Injection, effectively captures relation information and generates natural relations between customized objects while mitigating object confusion.
-We also provide a fair comparison with other methods on a new benchmark, RelationBench.
-The extensive experiments on three benchmarks demonstrate the effectiveness of our method, highlighting its potential for interactive scenario generation, relation detection, etc.
-6 Appendix
-This appendix provides additional details on the implementation of the Local Image Encoder and baseline methods.
-In the second section, we offer a brief introduction to the RelationBench benchmark.
-We then present further experimental results to validate the effectiveness of our method across various scenarios.
-Following that, we provide an in-depth analysis of the proposed components’ contributions to overall performance.
-Lastly, we discuss the potential ethical, privacy, and fairness implications of this technology, highlighting its social responsibility in real-world applications.
-6.1 More Implementation Details
+5, our model achieves the highest performance in the DINO score and ranks second in both the CLIP-T and CLIP-I metrics.
+8 Additional Implementation Details
 Local Image Encoder’s Implementation Details.
 To enhance region-language alignment of dense features, we employ self-distillation on CLIP-ViT-bigG.
-The training is conducted on the train2017 split of the COCO dataset for six epochs, using 8 A100 GPUs with a batch size of 2 per GPU.
+The training is conducted on the train2017 split of the COCO dataset for 6 epochs, using 8 A100 GPUs with a batch size of 2 per GPU.
 We apply the Adam optimizer with a learning rate of $51e-$ 5 and a weight decay of $0.1$.
-The Local Image Encoder, containing 1.8B parameters, extracts local tokens that are concatenated with image tokens during fine-tuning and inference to improve the alignment of object appearance.
-This distillation process enhances the performance of our base model, MS-Diffusion, in multi-object generation tasks.
-However, significant object confusion still occurs during relation generation.
+The Local Image Encoder, containing 1.8B parameters, extracts local tokens that are concatenated with image tokens during fine-tuning and inference to mitigate the confusion problem between objects.
 Baselines’ Implementation Details.
-For tuning-based methods, we implement Textual Inversion, DreamBooth, and Custom Diffusion using their respective Diffusers versions, with learning rates and tuning steps aligned to those reported in the original papers.
+We incorporate ReVersion [11] with MS-Diffusion [35] by its official implementation.
+We fine-tuning a learnable text embedding on a set of relation-specific images.
+We inject the text embedding into text prompt embedding during inference.
+We implement ReVersion [11] on DreamBooth [31] by similar procedure.
+For tuning-based methods, we implement Textual Inversion [7], DreamBooth [31], and Custom Diffusion [15] using their respective diffuser versions, with learning rates and tuning steps aligned to those reported in the original papers.
+We implement Mix-of-Show [8] by their official repository.
 We utilize the official implementations and checkpoints for training-based methods, adjusting hyperparameters as needed during evaluation.
-Specifically, we set the scale to 0.6 in MS-Diffusion and sample 30 steps using the EulerDiscreteScheduler.
-For the SSR Encoder, we employ the UniPCMultistepScheduler, sampling 30 steps and adjusting the scale for each object to accommodate different cases.
+Specifically, we set the scale to 0.6 in MS-Diffusion [35] and sample 30 steps using the EulerDiscreteScheduler.
+For the SSR Encoder [43], we employ the UniPCMultistepScheduler, sampling 30 steps and adjusting the scale for each object to accommodate different cases.
 For $\lambda$-Eclipse, we apply the default settings of the official implementation without modification.
-6.2 RelationBench
-In this section, we show the object and text prompt contained in our RelationBench in Fig. 10.
+9 Failure cases
+As shown in Fig.
+18, we present three typical failure cases from our experimental results.
+The first involves unreasonable relation generation requests, such as asking a plushie octopus, which lacks limbs, to ’shake hands.’ In response, our model generates additional arms for the plushie octopus, leading to a mismatched appearance.
+The second issue is the unnaturalness of some generated relations, such as a duck failing to make contact with a cat as it should.
+The final failure case is object confusion at the interaction point, which is a common challenge across all multi-object generation models.
+Figure 18: Failure cases of our DreamRelation.
+10 RelationBench
+In this section, we show the objects and text prompts contained in our RelationBench in Fig.
+19, Fig.
+20, and Tab.
+LABEL:tab:catpion.
+The objects are selected from well-known benchmarks DreamBench and CustomConcept101, covering the commonly seen categories in the real world.
+The relations in text prompts have covered the most common relations in the real world.
 Figure 19: Objects in our proposed RelationBench
@@
 No. Prompt
-1 A { } is playing guitar on a park bench, serenading passersby.
-2 A { } is playing piano in a grand hall filled with glowing chandeliers.
-3 A { } is eating dinner in a bustling restaurant with soft jazz playing in the background.
-4 A { } is dancing in the moonlight on a rooftop terrace with a stunning city view.
-5 A { } is lifting weights in a modern gym, pushing through an intense workout.
-6 A { } is reading a book by the fireplace on a rainy evening.
+1 A { } is playing guitar on a park bench.
+2 A { } is playing piano in a grand hall.
+3 A { } is eating dinner in a bustling restaurant.
+4 A { } is dancing in the moonlight.
+5 A { } is lifting weights in a modern gym.
+6 A { } is reading a book by the fireplace.
 7 A { } is skiing down a steep slope in the Alps, with snowflakes falling gently.
 8 A { } is sleeping peacefully in a hammock under the shade of a palm tree.
-9 A { } is cooking lunch in an open-air kitchen overlooking a lush green valley.
-10 A { } is singing on stage during a vibrant music festival with thousands of fans.
-11 A { } is riding a bike along the scenic countryside road during sunset.
-12 A { } is riding a horse through a dense forest, surrounded by nature.
-13 A { } is riding a motorbike on a winding mountain road, feeling the wind rush by.
-14 A { } is playing soccer on a sandy beach, with a vibrant sunset in the background.
-15 A { } is playing chess with a { } in a quiet park under the shade of a tree.
-16 A { } is partner dancing with a { } in a vintage ballroom with live music playing.
-17 A { } is carrying a { } across a rushing river, carefully finding their steps.
-18 A { } is fencing with a { } in an elegant arena, with spectators watching closely.
-19 A { } shakes hands with a { } at a formal awards ceremony after receiving a prize.
-20 A { } is kissing a { } in the rain under a large umbrella in a romantic city square.
+9 A { } is cooking lunch in a kitchen.
+10 A { } is singing on stage during a vibrant music festival.
+11 A { } is riding a bike along the scenic countryside road.
+12 A { } is riding a horse on the grassland.
+13 A { } is riding a motorbike on the street.
+14 A { } is playing soccer on football playground.
+15 A { } is playing chess with a { } under a tree.
+16 A { } is partner dancing with a { } in a vintage ballroom.
+17 A { } is carrying a { } on the diving room.
+18 A { } is fencing with a { } in an elegant arena.
+19 A { } shakes hands with a { } in the forest.
+20 A { } is kissing a { }.
 21 A { } is playing basketball with a { } on a street court.
@@
 25 A { } is sitting back to back with a { } on a hilltop.
-6.3 More Results of Relation-aware Image Customization
-More Results of Single-object relation-aware generation. We show more qualitative results in Fig. 12. We compare our methods with both training-based and tuning-based methods.
-Figure 12: Single-object comparison. TI and DB indicate Textual Inversion and DreamBooth, respectively. Our methods achieve the best balance between relation generation and identity preservation.
 Figure 21: Single-object comparison with our base model MS-Diffusion
-More Results of Multi-object relation-aware generation.
-In this section, we present additional results on multi-object relation-aware generation.
-In addition to implementing our RelationBooth on MS-Diffusion, we also use SDXL as a base model.
-As shown in Fig.
-15, our RelationBooth enhances SDXL’s ability to generate images that accurately depict object relations.
-Compared to ReVersion (Huang et al., 2023), our method generates more precise relations, without object confusion or missing objects, demonstrating the strong performance of RelationBooth in the Relation Generation task.
-Figure 14: Multi-object relation-aware image customization results of pet, toy, plushie, and person.
-Figure 15: Our RelationBooth is compatible with SDXL to address Relation Inversion task.
-Failure cases.
-As shown in Fig.
-16, we present three typical failure cases from our experimental results.
-The first involves unreasonable relation generation requests, such as asking a plushie octopus, which lacks limbs, to ’shake hands.’ In response, our model generates additional arms for the plushie octopus, leading to a mismatched appearance.
-The second issue is the unnaturalness of some generated relations, such as a duck failing to make contact with a cat as it should.
-The final failure case is object confusion at the interaction point, which is a common challenge across all multi-object generation models.
-Figure 16: Failure cases of our RelationBooth.
-6.4 Additional Analysis on Our Proposed Components
-Our well-curated Data.
-Our curated data are generated by DALLE-3 (Betker et al., 2023), which can effectively preserve the identity of common categories, such as animals, through its multi-turn dialogue capability.
-While our fine-tuning dataset consists solely of animals, our method focuses primarily on the relation information in the ground truth by disentangling identity using an off-the-shelf identity extractor.
-Therefore, the specific categories in the fine-tuning data are not the most critical factor for relation learning.
-Figure 17: Our fine-tuning dataset as an example.
-Keypoint Matching Loss.
-We use X-Pose as our keypoint detector due to its open-vocabulary detection capabilities.
-The keypoint matching loss (KML) facilitates relation generation by explicitly guiding the model’s pose manipulation, resulting in more accurate pose generation.
-As shown in Fig.
-19, the cat’s arm crosses the dog’s body, accurately depicting the ’hug’ relation.
-Figure 18: Illustration of Keypoint Matching Loss
-Figure 19: Additional ablation studies on KML and Local Tokens.
-Local Tokens Visualization.
-To understand why local features enhance relation-aware generation, we adopt Principal Component Analysis (PCA) to project the dense feature more compactly.
-As shown in Fig.
-20, dense features provide more fine-grained information than CLIP image tokens, aiding in distinguishing between different objects during the generation process and helping to avoid object confusion, especially in cases of heavy overlap.
-Moreover, it can facilitate object appearance alignment.
-Figure 20: Visualization of Dense feature and Image tokens by Principal Component Analysis (PCA).
-6.5 Social Compact
+11 Incorporate with CogVideoX-5b-I2V
+Additionally, we use our generated relation-aware customized images as the first frame to generate videos by the CogVideoX-I2V model, the generated results are shown in Fig.
+22.
+CogVideoX-I2V can handle simple relations such as ”shaking hands” but struggles with complex interactions such as ”hugging”.
+Figure 22: Incorporate with CogVideoX-5b-I2V.
+12 Social Impact
 Positive societal impacts.
@@
 Promoting transparency about generated content and integrating fairness and diversity considerations into dataset selection are key strategies for mitigating potential harms.
-6.6 Incorporate with CogVideoX-5b-I2V
-Figure 21: Incorporate with CogVideoX-5b-I2V.